A Complete Bibliography of The Bell System Technical Journal, 1930–1939

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

16 March 2020 Version 1.04

Title word cross-reference

n [Rio39].

1932 [OD32]. 1936 [Ano36i]. 1938 [Har39].

21-Type [Cri31].

Above [Fos33]. Absolute [Siv31]. Absorption [Sch38]. Abstracts [Ano30a, Ano30b, Ano30c, Ano30d, Ano31a, Ano31b, Ano31c, Ano31d, Ano32a, Ano32b, Ano32c, Ano32d, Ano33a, Ano33b, Ano33c, Ano33d, Ano34a, Ano34b, Ano34c, Ano34d, Ano35a, Ano35b, Ano35c, Ano35d, Ano36a, Ano36b, Ano36c, Ano36d, Ano37a, Ano37b, Ano37c, Ano37d, Ano38a, Ano38b, Ano38c, Ano38d, Ano39a, Ano39b, Ano39c, Ano39d]. Account [Gou34]. Acetate [AK32, WB32]. Acoustic [Fla32, Mas30]. Acoustical [Wen35]. Action [BH36b, Gou34]. Adaptation [BK34]. Addendum [Bur37a]. Advances [Dar30a, Dar30b, Dar30c, Dar31, Dar32a, Dar32b, Dar33a, Dar33b, Dar34a, Dar34b, Dar34c, Dar35, Dar36, Dar37, Dar39]. Affecting [HS31, Sou31]. Against [MW34]. Aid [MW34]. Aids [Nan32]. Air [Nan32, TJO35]. Airplane [SLT36]. Alloy [Bur36b]. Alloys [Elm36, HTSP35]. along [Sun37]. Alternating [Leg36, MM39]. America [Cla30a]. Among [Sco32]. Amplification [JL35]. Amplifier [Doh36, SS37, Sam37]. Amplifiers [Bla34, Chr32, Fay32, HD39]. Amplitude [Nor38, Wri38]. Analysis [CH36, CR38]. Analyzer [Mon38]. Antenna [FF37]. Antennas [Bru31, BBL35, Sou31]. Anti [Gib38]. Anti-sidetone [Gib38]. Apparatus [Ive31a, Lan30]. Application [BM31, CF37, CC33, Cha39, Fru32, Law35, Sch36, Sch38]. Applications [Leg39, Mol32, Nor37, SM39]. Applied [Col37b]. Applying [BRS39]. Approximate [Mas30]. Area [Han35]. Articles [Ano30a, Ano30b, Ano30c, Ano30d, Ano31a, Ano31b, Ano31c, Ano31d, Ano32a, Ano32b, Ano32c, Ano32d, Ano33a, Ano33b, Ano33c, Ano33d, Ano34a, Ano34b, Ano34c, Ano34d, Ano35a, Ano35b, Ano35c, Ano35d, Ano36a, Ano36b, Ano36c, Ano36d, Ano37a, Ano37b, Ano38a, Ano38b, Ano38c, Ano38d, Ano39a, Ano39b, Ano39c, Ano39d]. Articulation [CC33]. Artificial [Dar38]. Aspects [DH33, HO35, SS33]. Associated [BH36a]. Atmospheric [ECM38, Wal33]. Atmospherics [BB33]. Atoms [Dar37]. Attenuation [GLC36]. Audible [Sno31]. Audio [BB33, Gus38]. Auditoriums [Mac30]. Auditory [ACM34, BK34, Fle34, Scr34, SS34, WT34].

B [Fay32]. Background [Aik34]. Balanced [Cla35, MS37]. Band [Aff37, Cla35, ES34, Ham34, Mas37, MS37, MP37]. Band-Pass [Mas37]. Barrie [AG39]. Basic [Fle34]. Battery [Gib38, HTSP35]. Bayes [Mol31]. Bearing [Gou34]. Behavior [BH36b]. Bell [Ano30a, Ano30b, Ano30c, Ano30d, Ano31a, Ano31b, Ano31c, Ano31d, Ano32a, Ano32b, Ano32c, Ano32d, Ano33a, Ano33b, Ano33c, Ano33d, Ano34a, Ano34b, Ano34c, Ano34d, Ano35a, Ano35b, Ano35c, Ano35d, Ano36a, Ano36b, Ano36c, Ano36d, Ano36i, Ano37a, Ano37b, Ano37c, Ano37d, Ano38a, Ano38b, Ano38c, Ano38d, Ano39a, Ano39b, Ano39c, Ano39d, Bro36, OD32, Wat38]. Between [CA36, FSS38, HO35, SO37, Wal33]. Boston [GM35]. Bridge [Fer33, HE31, Leg36, Mea38]. Broad [Aff37, MP37]. Broad-Band [Aff37, MP37]. Broadcasting [Aik34, Cla30b, Gil31, Nel30].

C [Fay32]. Cable [Abr35, AGC32, AG39, Bur36b, CIK38, CG30, CO32, CK33, HE31, HB35, Lan38, She31, Str38, WTD38]. Cables [BRS39, Cla30a, GG38, Kem31, Law35, WF33]. Caesium [PK32]. Calcium [HTSP35]. Calculation [Ben33, FM33]. Calculus [Car30a, Car36c]. Calibration [Siv31]. Campbell [Ano35j, Jew35]. Carbon [CP36, Gou34, Jon31b]. Carrier [AGC32, Aik31, BRS39, Car39b, CIK38, CK33, FAM38, GG38, Har32, IHW39, KA39, Lan38, OBB39, SM39, WTD38]. Carrying [TJO35]. Cathode [Fru32, Joh32]. Cell [PK32]. Cellulose [AK32, WB32]. Centenary [Ano31i]. Ceramic [JS39]. Certain [Abr33, Sou31]. Cessation [BH36a]. Channel [Aik34, FAM38, HD39, IHW39, Ive31a, KA39, Lan38, OBB39]. Characteristic [NM38]. Characteristics [BHD35, Fle31, Gre30, MG34, Nor38]. Charge [FSS38, Rac38]. Chemistry [Wil30a]. Circuit [Bes33, Bes38, CF37, Car39a, CG30, LWF34]. Circuits [Abr33, Abr35, Aff37, BM31, BG30, Chr32, CO32, Cla35, CD37, GLC36, HO35, LP34, LP35, Mar30, Mar31, Mas37, ME33, MP37, Pet30, PH39, PP36, Wis31, WM32]. Circulating [Abr35]. Class [Fay32]. Classification [Fer33]. Clearance [EN39]. Closing [Ghe31]. Clutch [Wie39]. Coaxial [ES34, MS37, Sch34, SO37, Str38]. Cobalt [Elm36]. Coil [WT31]. Commercial [Fru32]. Common [CD37, Gib38, Gil31]. Common-Neutral [CD37]. Communication [CH36, CR38, Fon30, GL36, GJ30, Nan32]. Compandor [MW34]. Company [JS39]. Compensated [Mas37]. Composite [DG39]. Computation [DG39]. Concept [Sch38]. Conception [Dav32]. Condenser [HF32, Jon31b, Siv31]. Conductivity [MM39, Wal33]. Conductors [Sun37]. Considerations [Sou36]. Constant [Lle32, Nor37]. Constructing [Har32]. Construction [HTSP35, Jon31b]. Contemporary [Dar30a, Dar30b, Dar30c, Dar31, Dar32a, Dar32b, Dar33a, Dar33b, Dar34a, Dar34b, Dar34c, Dar35, Dar36, Dar37, Dar39]. Content [Wal33]. Continuously [Zin30]. Contributors [Ano30e, Ano30f, Ano30g, Ano30h, Ano31e, Ano31f, Ano31g, Ano31h, Ano32e, Ano32f, Ano32g, Ano32h, Ano33e, Ano33f, Ano33g, Ano33h, Ano34e, Ano34f, Ano34g, Ano34h, Ano35f, Ano35g, Ano35h, Ano35i, Ano36e, Ano36f, Ano36g, Ano36h, Ano37f, Ano37g, Ano37e, Ano37h, Ano38e, Ano38f, Ano38g, Ano38h, Ano39e, Ano39f, Ano39g, Ano39h]. Control [She30, Tay37, Wri38]. Controlling [Nor38]. Conversation [NM38]. Conversations [FCK30]. Cooperative [MH31]. Coordination [CD37, Ghe31, HS31, Pac31]. Copper [Car39b]. Corpuscle [Dar30a]. Correction [Bod30, Pay30]. Correlation [Ske36]. Corrosion [Bur36b, Bur36a, BH36b]. Cosmic [Dar32a, Dar39]. Cotton [Wal33]. Coupling [Chr32]. Covered [She31]. Craft [Han35]. Creosoted [CA36]. Crossbar [SR39]. Crosstalk [Cha34a, Cha34b, HB35, SO37, WTD38]. Crystal [LWF34, Lan38, Mas35, Mas37]. Crystals [Mas34, Sho39]. Current [HB35, Leg36, MM39]. Currents [Abr35, Sun37]. Curve [Haw31]. Curves [Bur37b, Bur37a]. Cycle [And39, Str37]. Cycles [Cur33]. Cylindrical [Sch34].

Data [Dar32a, ME33]. Decay [CA36]. Definition [FM33]. Demonstration [Dav32]. Densities [EL37, Leg36]. Design [MD30, Zob31]. Detection [Aik31]. Determination [BH36b]. Development [AK32, Bow37, CW31, IHW39, JI32, WB31, WB32]. Developments [Ano36i, Bru31, CC33, Cla30a, Fon30, Gil31, JS39, SCC39, She31]. Devices [Nor38]. Dial [SR39, Wie39]. Dielectric [MM37, MM38, MM39]. Differ [Aik31]. Digests [Ano35k, BBK36, Bow37, BH36b, Car36a, Doh36, FKSW36, GM35, Han35, HTSP35, Mas35, Sam37, SG36, Ske36, SLT36, Sun37, Wil35]. Diodes [LB39, Rac38]. Dipole [Ric37]. Direct [MM39]. Directional [MR36]. Directive [Bru31, Sou31]. Directivity [BB35]. Discoverer [Bla38]. Discoveries [OD32]. Discovery [Dav38]. Distance [CG30, Cla30a, CO32]. Distortion [KB35, Lan30, NB30, Ste30]. Distribution [CD37, HB35]. Disturbances [Car31]. Dr. [Ano35j, Jew35].

Ear [IGJ32]. Earth [Bur37a, Bur35, Bur37b, Car36a, Fos31, Fos33, Ric37, Rio31, RS33]. Echo [Hor38]. Eclipse [SG36]. Economic [She30]. Effect [Aik34, FKSW36, Hor38, HB35, Rac38, Wal33, Wis31]. Effective [ME33]. Effects [SG36, Ste30]. Efficiency [Abr33, Doh36]. Efficient [HF32]. Elastic [PL36]. Electric [Bla38, Bod35, CF37, FKSW36, Har32, JS39, Zob31]. Electrical [Ghe32, Mas34, OD32, Wal33]. Electrode [LP34, LP35, Pid35]. Electrodes [FSS38]. Electromagnetic [Sch34]. Electromagnetics [Sch36]. Electromechanical [HW36, Mas35]. Electron [Bec35, Dav32, Dav38, Gra39, Pet39]. Electron-Optics [Gra39]. Electronics [Lle34]. Electrons [Dar30c, Dar37, Sho39]. Electrostatic [Gra39]. Elements [LWF34, Mas34]. Emission [Bec35]. Employing [Mas34]. Energies [Sho39]. Energy [Car30b, Car31]. Energy-Frequency [Car31]. Engineering [Hoy33, WB32]. Equalizers [Bod38].

Equipment [Osw30]. Equivalence [Sch36]. Equivalent [Lle36, PH39]. Evaluation [WLH37]. Exchange [KLS36, PB36]. Expansion [Mol32, Mol36]. Experience [BRS39]. Experiment [PKW34]. Experimental [Sou36]. Experiments [BB35]. Explanation [Gib38]. Exponential [Bur38]. Expository [Mol31]. Extension [Car36c, LP34]. Extensions [EW31, LP35, Zob31]. Extraneous [TJO35].

Facilities [KLS36]. Factor [Bla32]. Factors [Abr33, Sou31, SS34]. Fading [BB35]. Faraday [Ano31i]. Faults [HE31]. Features [Ing38, Ive31b, WTD38]. Feedback [Bla34, Car39a, Cha39, Ste38]. Ferromagnetic [Boz36, KB35]. Field [Bur37a, Bur37b]. Filter [Cle34, Nor37]. Filters [Bod35, BD35, Har32, Lan38, Mas30, Mas34, Mas37, MS37, Pay30, Zob31]. Fine [Zin30]. First [Dar32b, Dar33b]. Flood [Har39]. Fluctuation [Pea34]. Fluctuations [CP36]. Flux [EL37, Leg36]. Forces [Car36b]. Fourth [Dar35]. Frequencies [Gus38, Lle34, Lle36, MS37, SS37, Sam37, TJO35]. Frequency [Aff37, Aik31, BB33, Car31, CMS36, CF37, Car39a, Cha39, Chr32, CW31, Dar32b, Dar33a, Gil31, GLC36, HO35, KB35, KS35, Lle32, Lle35, LB39, Mar30, Sno31, Sou36, WB31]. Frequency-Modulation [CF37, Car39a, Cha39]. Function [Ric37]. Functions [Mol32, Mol36]. Fundamentals [Wat38]. Further [ECM35, LP35]. Fusion [Dar30a]. Future [GJ32].

Gain [Sou31]. Gamma [Mol32]. Gases [Dar30b, Dar32b, Dar33a, SS33]. General [Bod35, Osb30, Sou36]. Generated [TJO35]. Generation [PMW37]. Geological [Car36a]. George [Jew35]. Graded [Wil31]. Granular [CP36]. Grid [Lle36, Sam37]. Ground [BG30, Sun37, Wis31]. Ground-Return [BG30, Sun37].

Grounded

[Fos31, Fos33, Pet30, Rio31, RS33]. Group [PMW37]. Groups [Nor37]. Guides [CMS36, Sou36].

Handset [JI32]. Harbor

[Han35, Wil35, GM35]. Harmonic [Mon38]. Harmonics [PMW37]. Havana [AGC32]. Heart [FKSW36]. Heaviside [Car30a]. Henry [Ghe32, OD32]. Hermitian [Mol36]. Hertz [Bla38]. High [Chr32, Dar32b, Dar33a, Doh36, GLC36, Her38, KS35, Lle35, Lle36, LB39, MS37, Mol36, SS37, Sam37, Lle34]. High-Efficiency [Doh36]. High-Frequency [Chr32, Dar32b, Dar33a, GLC36, KS35]. Horizontal [BBL35]. Horizontally [RS33]. Humidity [Wal33]. Hurricane [Har39]. Hyper [CMS36, Sou36]. Hyper-Frequency [CMS36, Sou36].

Ideal [BD35]. II [Bur36b]. III [MM39]. Image [IGB30]. Impedance [Bod30, Fla32, Fos31, Fos33, Pay30, Pet39, RS33, Sch38]. Impedances [BG30, Cri31, Fer33, HM35, Rio39]. Improved [OBB39]. Improvements [GL36, LWF34]. Incomplete [Mol32]. Indicator [EN39]. Induction [CW31, DH33, HO35, WB31]. Inductive [CD37]. Industry [Wil30a]. Infinite [Pet30]. Inhibitive [BH36b]. Instruments [Jon38, Wen35]. Insulated [She31]. Insulating [Kem31, MM37, MM38, MM39]. Insulation [AK32, WF33, WB32]. Insulators [Wil30b]. Integrals [Mol32]. Intense [TJO35]. Interception [Dar30c]. Interconnection [Wil31]. Intervals [NM38]. Ionization [Dar30b]. Ions [Dar30c]. Ionsphere [SG36]. Iron [BH36b, Elm36]. Irregularities [ECM38, MP37]. Issue [Ano30e, Ano30f, Ano30g, Ano30h, Ano31e, Ano31f, Ano31g, Ano31h, Ano32e, Ano32f, Ano32g, Ano32h, Ano33e, Ano33f, Ano33g, Ano33h, Ano34e, Ano34f, Ano34g, Ano34h, Ano35f, Ano35g, Ano35h, Ano35i, Ano36e, Ano36f, Ano36g, Ano36h, Ano37f, Ano37g, Ano37e, Ano37h, Ano38e, Ano38f, Ano38g, Ano38h, Ano39e, Ano39f, Ano39g, Ano39h].

J [SM39]. Joint [CW31, MH31, WB31]. Joseph [OD32]. Journal [Ano36i, OD32]. July [Ano36i, OD32].

Key [AGC32].

Laboratory [WLH37]. Ladder [Rio37]. Laplace [Mol36]. Laplacian [Mol32, Mol36]. Larynx [Weg30]. Layer [RS33]. Lead [Bur36b, HTSP35, She31]. Lead-Alloy [Bur36b]. Lead-Calcium [HTSP35]. Lead-Covered [She31]. Leaky [Sun37]. Length [Pet30]. Light [Dar30b]. Limiting [Abr33]. Limits [JL35]. Line [Bur38, CH38, Cla30b, Cle34, IHW39, Wil30b]. linear [Har36]. Lines [ACM34, ES34, Gre30, KA39, MS37, Pil39, Sch34, SO37, SF32]. Load [HD39]. Loaded [Zin30]. Locating [HE31]. Locations [WM30]. Lockout [Hor38]. Long [BM31, BT35, BH36a, CG30, Cla30a, CO32]. Long-Wave [BM31, BT35, BH36a]. Losses [EL37]. Loud [WT34]. Loudness [FM33]. Low [CW31, EL37, HO35, Leg36]. Low-Frequency [CW31, HO35]. Lying [Fos31, Fos33].

Magnet [Sco32]. Magnetic [ES35, EL37, Elm36, Gus38, Haw31, Hic37, Leg36, Leg39, PMW37, Sco32]. Magnetism [Dar36]. Magneto [PL36]. Magneto-Elastic [PL36]. Magnets [Sco32]. Maintenance [Bes38]. Manufacture [She31]. Manufactured [She30]. Manufacturing [AK32, JS39]. Marine [GM35]. Material [Kem31]. Materials [ES35, Fon30, Leg39, MM37, MM38, MM39, TC39]. Mathematical [CMS36]. Means [LB39]. Measurement [BBK36, FM33, NB30, Pen34, SCC39]. Measurements [IGJ32, Leg36]. Measuring [Bes33, Bes38, Fer33, Fla32]. Mechanism [Bur36a]. Memoranda [Ano35j]. Message [Mar30]. Metals [Bur36b, Bur36a, SS33]. Method [Bod30, Fla32]. Methods [CH36, Fer33, Har32, HE31]. Microchemical [CH36]. Microphone [Gou34, HF32, MR36]. Microphones [CP36, Jon31b, WT31, WT34]. Microphonic [Pen34]. Million [Str37]. Million-Cycle [Str37]. Miniature [HF32]. Minimum [GLC36]. Mobile [BHD35]. Modulated [Aik31, Doh36]. Modulation [Ben33, CF37, Car39a, Cha39]. Modulator [PH39]. Modulators [Car39b]. Moisture [Wal33, Wal37]. Motion [Her38, Qua30]. Moving [WT31]. Moving-Coil [WT31]. Multi [HD39, Ive31a, LP35, Pid35]. Multi-Channel [HD39, Ive31a]. Multi-Electrode [LP35, Pid35]. Multiple [FF37]. Multiples [Wil31]. Music [ACM34, BK34, Fle31, Fle34, Sno31, SS34, WT34]. Mutual [BG30, Fos31, Fos33, HM35, RS33].

National [Cla30b]. Natural [Dar38]. Nature [Dar32a, Luc30]. Negative [Cha39, Cri31, Lle36, Sam37]. Negative-Grid [Lle36, Sam37]. Network [Rio37, Rio39]. Networks [EW31, Lle36, Mas30]. Neutral [CD37]. Nickel [Elm36]. Noise [Aik34, BBK36, DH33, DG39, Pea34, Pen34, PL36, Rac38, Sno31, SLT36, Tay37, WTD38, WM30, WB31]. Non [Har36, HB35, MR36]. Non-Directional [MR36]. Non-linear [Har36]. Non-Uniform [HB35]. Notes [Car30a]. Nucleus [Dar33b, Dar34a, Dar34b, Dar34c, Dar35]. Number [Law35, Rio39].

Oblique [Car36b]. Occasioned [Hor38].

Occurrence [Hor38]. One [Pet30]. Open [Cha34a, Cha34b, Gre30, Ham34, IHW39, KA39]. Open-Wire [Cha34a, Cha34b, Gre30, Ham34, IHW39, KA39]. Operation [Bro36, Fay32, Lle35]. Operational [Car30a, Car36c]. Optical [Ive31b, Mon38]. Optics [Gra39]. Optimum [Mac30]. Order [Mol36]. Oscillations [Har36, HW36, LB39]. Oscillator [Mea38, Sam37]. Oscillators [KS35, Lle32, Ste38]. Oscillograph [Cur33, Joh32]. Other [CP36]. Overseas [EW31]. Oxide [Car39b]. Oxygen [PK32].

Painted [BH36b]. Paints [BH36b]. Paper [She31]. Paper-Insulated [She31]. Papers [Ano37c, Ano37d]. Paragutta [Kem31]. Parallel [FSS38, HM35, Pet30]. Part [Dar32b, Dar33a, Dar33b, Dar34a, Dar34b, Dar34c, Dar35]. Particles [Dar39]. Pass [Mas37]. Penetration [CA36]. Performance [Mar31]. Permalloy [EL37]. Permanent [Sco32]. Permeability [Wis31]. Perspective [ACM34, BK34, Fle34, Scr34, SS34, WT34]. Phase [Lan30, NB30, Ste30]. Phenomena [BT35, BH36a, Dar32b, Dar33a, ECM33, ECM35, Nel30, Ske36]. Photoelectric [PK32]. Photography [Her38]. Physical [Fle31, MPS32, SS34, Wis37]. Physics [Dar30a, Dar30b, Dar30c, Dar31, Dar32a, Dar32b, Dar33a, Dar33b, Dar34a, Dar34b, Dar34c, Dar35, Dar36, Dar37, Dar39, Sho39]. Picture [Her38]. Piezoelectric [Mas35]. Pine [CA36]. Pioneer [Ghe32]. Plan [Osb30]. Plane [Bur37a, Bur37b, FSS38]. Plant [Ghe31, Pac31]. Plastic [TC39]. Poles [CA36, Jon31a, MH31]. Potentials [Sun37]. Power [BBK36, CD37, Doh36, Ghe31, HS31, HO35, Pac31, SS37, Sch38]. Practical [DH33, SS33]. Practise [HS31]. Precision [Har32]. Present [Boz36]. Presentation [Mol31]. Preservatives [WLH37]. Printing [BM31]. Probability [Hoy33]. Problems [CH38, GJ32, IHW39, Sch36, Sch38]. Process [AK32]. Processes [Bur36a]. Product [She30]. Production [LB39]. Products [Ben33]. Program [CG30, Cle34, Ham34]. Propagation [Bur35, BHD35, Bur37a, Bur37b, SBF33, Zin30]. Properties [MPS32, MM37, MM38, MM39, Pet39, Sco32]. Proportioning [GLC36]. Puget [Han35]. Pulp [WF33].

Quality [She30, Ste30]. Quantum [Sho39]. Quartz [LWF34, Mas34].

Radiating [Ric37]. Radiation [Sch36]. Radio [BM31, BT35, BH36a, Bow37, Bur35, Bur37b, EW31, GM35, Han35, MS37, MW34, Nel30, PP36, SP35, Ske36, SF32, Tay37, WE30, WM32, Bur37a]. Radioactivity [Dar38]. Random [Car31, DG39]. Range [Aff37, Mar30, Wri38]. Ranges [Sno31]. Rating [HD39, Mar31, ME33]. Ray [Joh32]. Rays [BH36a, Dar32a, Dar39]. Reactance [Har36]. Reality [Wis37]. Receiver [Col37b, Tay37]. Receivers [MD30, WT31]. Receiving [Car39a]. Reception [FF37]. Reciprocal [Car30b]. Recording [Bes33, Hic37]. Reduction [BB35, Pen34, SLT36, Tay37]. Reflection [Sch38]. Refraction [Sch38]. Regeneration [Nyq32, PKW34]. Related [BT35, Nel30]. Relation [CA36, ES35, MG34, Wal33]. Relationships [Sco32]. Reliability [PP36]. Remarks [Ghe31]. Repeater [Cri31]. Repeatered [Abr33]. Representation [And39, Mas35]. Reproduction [ACM34, BK34, Fle34, Scr34, SS34, WT34]. Requirements [Fle34]. Research [CH36, CR38, Col37b, CW31, WB31]. Researches [Gou34]. Resistance [CP36, HE31, Mas37, Nor37]. Resistances [CP36]. Resistivity [Car36a]. Resulting [DG39]. Results [Ben33, ECM33, ECM35, Sou36]. Return [BG30, Sun37, Wis31]. Reverberation [Mac30]. Rhombic [BBL35]. Room [WM30].

Scanning [MG34]. Scattering [Dar30c]. Scientific [Col37b]. Sea [WE30]. Second [Dar33a, Dar34a]. September [Har39]. Series [Ric37]. Service [Bow30, GM35, Han35, Osb30, Osw30, PB36, Wil35, WE30]. Services [Ano36i]. Set [Gib38]. Sets [Ing38, Jon38, Wil35]. Shape [BBK36]. Shared [Aik34]. Sheathing [Bur36b]. Sheet [EL37]. Shielded [GLC36]. Shielding [Gus38, Sch38]. Shields [Sch34]. Ship [Han35, Wil35]. Ship-to-Shore [Han35, Wil35]. Ships [WE30]. Shock [FKSW36]. Shore [Han35, Wil35]. Short [Bow30, Bru31, BHD35, ECM33, ECM35, ECM38, FF37, Osw30, PS35, PP36, SBF33, SP35, SF32]. Short-Wave [Bow30, Bru31, FF37, Osw30, PS35, PP36, SF32]. Shot [Rac38]. Sideband [PS35]. sidetone [Gib38]. Signal [MG34]. Signaling [KLS36]. Signals [Nor38]. Silver [PK32]. Singing [Abr35]. Single [FAM38, PS35]. Single-Sideband [PS35]. Slightly [Aik31]. Solar [Ske36]. Solders [MPS32]. Solids [Sho39]. Some [Cla30a, DH33, ECM33, Fle31, Gil31, Gou34, HO35, Ive31b, JS39, LWF34, MPS32, Mol32, SS33, Sch36, SM39]. Sound [BB30, Han35, Hic37, TJO35]. Sounds [FCK30]. Source [PL36]. Sources [Ano30a, Ano30b, Ano30c, Ano30d, Ano31a, Ano31b, Ano31c, Ano31d, Ano32a, Ano32b, Ano32c, Ano32d, Ano33a, Ano33b, Ano33c, Ano33d, Ano34a, Ano34b, Ano34c, Ano34d, Ano35a, Ano35b, Ano35c, Ano35d, Ano36a, Ano36b, Ano36c, Ano36d, Ano37a, Ano37b, Ano37c, Ano37d, Ano38a, Ano38b, Ano38c, Ano38d, Ano39a, Ano39b, Ano39c, Ano39d, DG39]. Southern [CA36]. Space [FSS38, Rac38]. Spacing [CH38]. Speakers [WT34]. Special [CH36]. Specifications [Jon31a]. Spectrochemical [CR38].

Spectrum [Car31]. Speech [Fle31, Sno31]. Speed [Her38]. Spherical [Bur35]. Spinning [Dar37]. Splicing [Law35]. Spontaneous [CP36]. Spring [Wie39]. Sputtering [Fru32]. Stabilized [Bla34, Mea38, Ste38]. Standard [Jon31a]. States [Cla30a, Col37a, GJ30]. Static [MW34]. Stations [JI32]. Statistical [Car31]. Status [Boz36, CW31, MH31, WB31]. Steel [PL36]. Steels [Sco32]. Steerable [FF37]. Steering [BB35]. Storage [HTSP35]. Stratified [RS33]. Streams [Pet39]. Strength [Bur37a, Bur37b]. Structure [Car36a, ES35, Luc30]. Studies [CH38]. Study [EL37, ECM33, ECM35, Wil30b]. Submarine [Kem31]. Subscriber [Gib38]. Sun [BH36a]. Sunspot [And39]. Superiorities [HTSP35]. Supplement [Ano36i, OD32]. Suppressors [Hor38]. Surface [Fos31, Fos33, Ric37, Rio31, Wis37]. Survey [ES35, Leg39, WM30]. Switchboards [KLS36]. Switching [Osb30, SR39]. Symphonic [ACM34, BK34, Fle34, SS34, WT34]. Symposium [ACM34, BK34, Fle34, Ghe31, Pac31, SS34, WT34]. Synchronization [Sto30]. System [Ano36i, BK34, Bes33, BB30, Cle34, FAM38, GJ30, GG38, Ham34, HF32, HW36, IHW39, IGB30, KA39, KLS36, Lan38, Leg39, ME33, OBB39, OD32, PB36, PS35, Rey36, SP35, SR39, SM39, Sto30, Str37, Str38, WTD38, Ano30a, Ano30b, Ano30c, Ano30d, Ano31a, Ano31b, Ano31c, Ano31d, Ano32a, Ano32b, Ano32c, Ano32d, Ano33a, Ano33b, Ano33c, Ano33d, Ano34a, Ano34b, Ano34c, Ano34d, Ano35a, Ano35b, Ano35c, Ano35d, Ano36a, Ano36b, Ano36c, Ano36d, Ano36i, Ano37a, Ano37b, Ano37c, Ano37d, Ano38a, Ano38b, Ano38c, Ano38d, Ano39a, Ano39b, Ano39c, Ano39d, Bro36, Wat38]. Systems [Aff37, BRS39, Bes38, Car39b, Cha39, Cla30b, CD37, ES34, Har32, Har36, SF32, Wil31].

Tandem [Bro36]. Tape [Hic37]. Technical [Ano30a, Ano30b, Ano30c, Ano30d, Ano31a, Ano31b, Ano31c, Ano31d, Ano32a, Ano32b, Ano32c, Ano32d, Ano33a, Ano33b, Ano33c, Ano33d, Ano34a, Ano34b, Ano34c, Ano34d, Ano35b, Ano35c, Ano35d, Ano35k, Ano36a, Ano36b, Ano36c, Ano36d, Ano36i, Ano37a, Ano37b, Ano37c, Ano37d, Ano38a, Ano38b, Ano38c, Ano38d, Ano39a, Ano39b, Ano39c, Ano39d, BBK36, Bow37, BH36b, Car36a, Doh36, FKSW36, GM35, Han35, HTSP35, Mas35, OD32, Sam37, SG36, Ske36, SLT36, Sun37, Wil35]. Telegraph [BM31, SCC39]. Telephone [Abr33, AGC32, Ano35e, BT35, BBK36, BRS39, Bes33, Bes38, Bla32, Bow30, Bow37, Car36b, CH38, Car39b, CIK38, CO32, CD37, Col37b, EW31, FAM38, FCK30, GJ30, Ghe31, GM35, Gre30, GG38, Han35, HS31, Hoy33, HO35, IGJ32, Ing38, JI32, Jon38, KA39, Lan30, Law35, Leg39, MD30, Mar30, Mar31, ME33, OBB39, Osb30, Osw30, Pac31, PL36, Pil39, PP36, Qua30, Qua36, SP35, SR39, She31, Ste30, Str37, Tay37, TC39, WF33, WTD38, WT31, Wil31, Wil30a, WM30, WE30, Wil30b, WM32]. Telephonic [NM38, Nor38]. Telephony [GJ32, MW34, PS35]. Telephotograph [Rey36]. Telephotography [MG34]. Teletypewriter [KLS36, PB36]. Teletypewriters [Wat38]. Television [BB30, IGB30, Ive31a, Ive31b, MG34, Sto30, Str38]. Temperature [Wal33]. Ten [Cur33]. Terminal [Rio39]. Terminals [CIK38]. Terms [Mol32]. Terrain [EN39]. Testing [Bes33, CC33]. Textile [AK32, WB32]. Textiles [Wal37]. Their [Jon31b, Sch36, Sco32]. Theorem [Car30b, Mol31, Rio37]. Theorems [Sch36]. Theoretical [DH33, SS33]. Theories [Dar30a]. Theory [Bod35, Boz36, CMS36, CF37, Car39a, Dar36, FSS38, HD39, Hoy33, Law35, LP34, LP35, MG34, Nyq32, PKW34, Pid35, Sch34, Weg30, Zob31]. Thermionic [Bec35]. Third [Dar34b, Dar34c].

Thousand [Cur33]. Three [LP34, OBB39]. Three-Channel [OBB39]. Three-Electrode [LP34]. Time [Bla32, Mac30, NM38, Rac38]. Toll [Ano36i, AG39, BRS39, Col37a, GG38, Osb30]. Toronto [AG39]. Tracer [Haw31]. Transatlantic [BT35, PS35]. Transcontinental [Pil39]. Transducer [Mas35]. Transformer [Chr32]. Transformers [GL36, Gus38, MS37]. Transients [Pet30, Rio31]. Transit [Rac38]. Transmission [ACM34, BT35, BH36a, BK34, Bes33, Bes38, Bla32, BB30, Bow30, BHD35, Bur38, CG30, Cla35, Col37a, ECM33, ECM35, ECM38, ES34, Fle34, Gre30, Hoy33, Ing38, IGB30, Mar31, MS37, ME33, MP37, Nel30, PB36, Sch34, SO37, SCC39, Ske36, SS34, SF32, Str38, WT34]. Transmitted [Aff37, Mar30, MG34]. Transmitter [Col37b]. Transmitters [MD30, Nel30, Siv31]. Transmutation [Dar31]. Transoceanic [Bow30, Bow37, Osw30]. Transportation [Nan32]. Treatment [AK32, WB32]. Trend [MD30]. Trends [Col37a, HS31]. Triode [Sam37]. Troostite [Luc30]. Tube [LP34, LP35, Lle34]. Tubes [Fay32, KS35, Lle35, Lle36, Pea34, Pen34, Pid35]. Twelve [IHW39, KA39]. Twelve-Channel [IHW39, KA39]. Twin [Cri31]. Two [Abr35, Aik31, BB30, Hor38, IGB30, Ive31b, KB35, RS33, Sto30, WM32]. Two-Frequency [KB35]. Two-Layer [RS33]. Two-Way [BB30, IGB30, Ive31b, Sto30, WM32]. Two-Wire [Abr35]. Type [Cri31, Qua36, SM39, Wie39].

Ultra [BHD35, ECM33, ECM35, ECM38, Lle34, Lle35, Lle36, LB39, SS37, Sam37, SBF33, SP35]. Ultra-High [Lle36, SS37, Sam37, Lle34]. Ultra-High-Frequency [Lle35, LB39]. Ultra-Short [SBF33]. Ultra-Short-Wave [BHD35, ECM33, ECM35, ECM38, SP35]. Unattended [SP35]. Unbalanced [Mas37]. Underground [Qua36]. Underlying [Ano36i]. Uniform [HB35]. Unit [FF37]. United [Cla30a, Col37a, GJ30]. Urban [BHD35]. Use [Jon31b, MH31, Mas37, MS37, TC39]. Used [Har32, Mas35, Wat38]. Using [Leg36].

Vacuum [Fay32, KS35, LP34, LP35, Lle34, Lle35, Lle36, Pea34, Pen34, Pid35]. Variable [Bod38, CF37, DG39]. Vibration [SLT36, Weg30]. Vodas [Wri37]. Voice [IGJ32, Tay37]. Volume [Abr33].

Wave [BM31, BT35, BH36a, BBK36, Bod35, BD35, Bow30, Bru31, BHD35, CMS36, Dar30a, ECM33, ECM35, ECM38, FF37, Har32, KB35, Mas34, Osw30, Pay30, PS35, PP36, Ric37, SBF33, SP35, Sou36, SF32, Wis37, Zin30, Zob31]. Wave-Filters [Zob31]. Waves [Aik31, Bla38, Dav32, Dav38, Doh36, TJO35]. Way [BB30, IGB30, Ive31b, Sto30, WM32]. West [AGC32]. Western [JS39]. Which [Aik31, Pet30]. Wide [Cla35, ES34, GJ32, Ham34, MS37]. Wide-Band [ES34, Ham34]. Wind [Qua30]. Winds [Car36b]. Wiping [MPS32]. Wire [Abr35, ACM34, BK34, CH38, Cha34a, Cha34b, Cla30b, EW31, Fle34, Gre30, Ham34, IHW39, KA39, MP37, Nan32, Qua36, SS34, WT34]. Wires [Car36b, Fos31, Fos33, HE31, HM35, HB35, PL36, Qua30, Rio31, RS33, Zin30]. Wood [Jon31a, WLH37]. Words [FCK30]. Work [MH31]. World [Ano35e, GJ32]. World-Wide [GJ32].

XIX [Dar30a]. XX [Dar30b]. XXI [Dar30c]. XXII [Dar31]. XXIII [Dar32a]. XXIV [Dar32b]. XXIX [Dar35]. XXV [Dar33a]. XXVI [Dar33b]. XXVII [Dar34a]. XXVIII [Dar34b, Dar34c]. XXX [Dar36].

XXXI [Dar37]. XXXII [Dar39].

Zenneck [Wis37].

References

Abraham:1933:CFL

[Abr33] Leonard Gladstone Abraham. Certain factors limiting the volume efficiency of repeatered telephone circuits. The Bell System Technical Journal, 12(4): 517–532, October 1933. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                                  tronic). URL http://bstj.bell-
                                                  labs.com/BSTJ/images/Vol12/
                                                  bstj12-4-517.pdf;   http://
                                                  www.alcatel-lucent.com/bstj/
                                                  vol12-1933/articles/bstj12-
                                                  4-517.pdf.

Abraham:1935:CCS

[Abr35] Leonard Gladstone Abraham. Circulating currents and singing on two-wire cable circuits. The Bell System Technical Journal, 14(4): 600–631, October 1935. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                                  tronic). URL http://bstj.bell-
                                                  labs.com/BSTJ/images/Vol14/
                                                  bstj14-4-600.pdf;   http://
                                                  www.alcatel-lucent.com/bstj/
                                                  vol14-1935/articles/bstj14-
                                                  4-600.pdf.

Affel:1934:TLS

[ACM34] H. A. Affel, R. W. Chesnut, and R. H. Mills. Transmission lines, symposium on wire transmission of symphonic music and its reproduction in auditory perspective. The Bell System Technical Journal, 13(2):285–300, April 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-2-285.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
2-285.pdf.

Affel:1937:TFR

[Aff37] H. A. Affel. Transmitted frequency range for circuits in broad-band systems. The Bell System Technical Journal, 16(4):487–492, October 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-4-487.pdf;   http://
www.alcatel-lucent.com/bstj/
vol16-1937/articles/bstj16-
4-487.pdf.

Aykroyd:1939:TBT

[AG39] M. J. Aykroyd and D. G. Geiger.

The Toronto–Barrie toll cable. The Bell System Technical Journal, 18(4):588–604, October 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-4-588.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
4-588.pdf.

Affel:1932:NKW

[AGC32] H. A. Affel, W. S. Gorton, and R. W. Chesnut. A new Key West–Havana carrier telephone cable. The Bell System Technical Journal, 11(2):197–212, April 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol11/
                                        bstj11-2-197.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol11-1932/articles/bstj11-
                                        2-197.pdf.

Aiken:1931:DTM

[Aik31] Charles B. Aiken. The detection of two modulated waves which differ slightly in carrier frequency. The Bell System Technical Journal, 10(1):1–19, January 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-1-1.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol10-
                                        1931/articles/bstj10-1-1.pdf.

Aiken:1934:EBN

[Aik34] C. B. Aiken. The effect of background noise in shared channel broadcasting. The Bell System Technical Journal, 13 (3):333–350, July 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol13/
                                        bstj13-3-333.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol13-1934/articles/bstj13-
                                        3-333.pdf.

Avery:1932:CAT

[AK32] C. R. Avery and H. Kress. Cellulose acetate treatment for textile insulation — development of the manufacturing process. The Bell System Technical Journal, 11 (2):231–244, April 1932. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-2-231.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
2-231.pdf.

Anderson:1939:RSC

[And39] C. N. Anderson. A representation of the sunspot cycle. The Bell System Technical Journal, 18 (2):292–299, April 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-2-292.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
2-292.pdf.

Anonymous:1930:ATAa

[Ano30a] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 9(1): 228–233, January 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
1-228.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-1-228.pdf.

Anonymous:1930:ATAb

[Ano30b] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 9(2): 398–403, April 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        2-398.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-2-398.pdf.

Anonymous:1930:ATAc

[Ano30c] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 9(3): 595–599, July 1930. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        3-595.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-3-595.pdf.

Anonymous:1930:ATAd

[Ano30d] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 9(4): 836–839, October 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        4-836.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-4-836.pdf.

Anonymous:1930:CIa

[Ano30e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 9(1): 234–235, January 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
1-234.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-1-234.pdf.

Anonymous:1930:CIb

[Ano30f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 9(2): 404–405, April 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
2-404.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-2-404.pdf.

Anonymous:1930:CIc

[Ano30g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 9(3): 600–602, July 1930. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
3-600.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-3-600.pdf.

Anonymous:1930:CId

[Ano30h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 9(4):

840–842, October 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
4-840.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-4-840.pdf.

Anonymous:1931:ATAa

[Ano31a] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 10(1): 149–152, January 1931. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-1-149.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
1-149.pdf.

Anonymous:1931:ATAb

[Ano31b] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 10(2):342–345, April 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-2-342.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
2-342.pdf.

Anonymous:1931:ATAc

[Ano31c] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 10(3):525–528, July 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-3-525.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
3-525.pdf.

Anonymous:1931:ATAd

[Ano31d] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 10(4):684–690, October 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-4-684.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
4-684.pdf.

Anonymous:1931:CIa

[Ano31e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 10(1):153–154, January 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-1-153.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
1-153.pdf.

Anonymous:1931:CIb

[Ano31f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 10(2):346–348, April 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-2-346.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        2-346.pdf.

Anonymous:1931:CIc

[Ano31g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 10(3):529–530, July 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-3-529.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        3-529.pdf.

Anonymous:1931:CId

[Ano31h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 10(4):691–692, October 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-4-691.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        4-691.pdf.

Anonymous:1931:FC

[Ano31i] Anonymous. The Faraday Centenary. The Bell System Technical Journal, 10(4):iii–vii, October 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-4-iii.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol10-1931/articles/bstj10-
4-iii.pdf.

Anonymous:1932:ATAa

[Ano32a] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 11(1): 185–194, January 1932. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-1-185.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
1-185.pdf.

Anonymous:1932:ATAb

[Ano32b] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 11(2):318–323, April 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-2-318.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
2-318.pdf.

Anonymous:1932:ATAc

[Ano32c] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 11(3):477–481, July 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-3-477.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol11-1932/articles/bstj11-
                                        3-477.pdf.

Anonymous:1932:ATAd

[Ano32d] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 11(4):622–628, October 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol11/
                                        bstj11-4-622.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol11-1932/articles/bstj11-
                                        4-622.pdf.

Anonymous:1932:CIa

[Ano32e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 11(1):195–196, January 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol11/
                                        bstj11-1-195.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol11-1932/articles/bstj11-
                                        1-195.pdf.

Anonymous:1932:CIb

[Ano32f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 11(2):324–325, April 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol11/
                                        bstj11-2-324.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol11-1932/articles/bstj11-
                                        2-324.pdf.

Anonymous:1932:CIc

[Ano32g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 11(3):482–483, July 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-3-482.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
3-482.pdf.

Anonymous:1932:CId

[Ano32h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 11(4):629–630, October 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-4-629.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
4-629.pdf.

Anonymous:1933:ATAa

[Ano33a] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 12(1): 119–122, January 1933. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-1-119.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
1-119.pdf.

Anonymous:1933:ATAb

[Ano33b] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 12(2):244–247, April 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-2-244.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
2-244.pdf.

Anonymous:1933:ATAc

[Ano33c] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 12(3):371–374, July 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-3-371.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
3-371.pdf.

Anonymous:1933:ATAd

[Ano33d] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 12(4):533–537, October 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-4-533.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
4-533.pdf.

Anonymous:1933:CIa

[Ano33e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 12(1):123–124, January 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-1-123.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
1-123.pdf.

Anonymous:1933:CIb

[Ano33f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 12(2):248–249, April 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-2-248.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
2-248.pdf.

Anonymous:1933:CIc

[Ano33g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 12(3):375–376, July 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-3-375.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
3-375.pdf.

Anonymous:1933:CId

[Ano33h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 12(4):538–539, October 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-4-538.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
4-538.pdf.

Anonymous:1934:ATAa

[Ano34a] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 13(1): 159–160, January 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-1-159.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
1-159.pdf.

Anonymous:1934:ATAb

[Ano34b] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 13(2):309–312, April 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-2-309.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
2-309.pdf.

Anonymous:1934:ATAc

[Ano34c] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 13(3):516–519, July 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-3-516.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
3-516.pdf.

Anonymous:1934:ATAd

[Ano34d] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 13(4):701–703, October 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-4-701.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
4-701.pdf.

Anonymous:1934:CIa

[Ano34e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 13(1):161, January 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-1-161.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
1-161.pdf.

Anonymous:1934:CIb

[Ano34f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 13(2):313–314, April 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol13/
                                        bstj13-2-313.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol13-1934/articles/bstj13-
                                        2-313.pdf.

Anonymous:1934:CIc

[Ano34g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 13(3):520–521, July 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol13/
                                        bstj13-3-520.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol13-1934/articles/bstj13-
                                        3-520.pdf.

Anonymous:1934:CId

[Ano34h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 13(4):704–705, October 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol13/
                                        bstj13-4-704.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol13-1934/articles/bstj13-
                                        4-704.pdf.

Anonymous:1935:AAB

[Ano35a] Anonymous. Abstracts of articles from Bell System sources. The Bell System Technical Journal, 14(1):173–175, January 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-1-173.pdf;   http://

www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
1-173.pdf.

Anonymous:1935:ATAa

[Ano35b] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 14(2):360–365, April 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-2-360.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
2-360.pdf.

Anonymous:1935:ATAb

[Ano35c] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 14(3):544–549, July 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-3-544.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
3-544.pdf.

Anonymous:1935:ATAc

[Ano35d] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 14(4):724–727, October 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-4-724.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol14-1935/articles/bstj14-
                                        4-724.pdf.

Anonymous:1935:AWT

[Ano35e] Anonymous. Around the world by telephone. The Bell System Technical Journal, 14(3):542–543, July 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-3-542.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        3-542.pdf.

Anonymous:1935:CIa

[Ano35f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 14(1):176–178, January 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-1-176.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        1-176.pdf.

Anonymous:1935:CIb

[Ano35g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 14(2):366–367, April 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-2-366.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        2-366.pdf.

Anonymous:1935:CIc

[Ano35h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 14(3):550–551, July 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-3-550.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
3-550.pdf.

Anonymous:1935:CId

[Ano35i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 14(4):728–730, October 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-4-728.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
4-728.pdf.

Anonymous:1935:DCM

[Ano35j] Anonymous. Dr. Campbell’s memoranda of 1907 and 1912. The Bell System Technical Journal, 14 (4):558–572, October 1935. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-4-558.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
4-558.pdf.

Anonymous:1935:TD

[Ano35k] Anonymous. Technical digests.

The Bell System Technical Journal, 14(4):698, October 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-4-698.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
4-698.pdf.

Anonymous:1936:ATAa

[Ano36a] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 15(1): 172–176, January 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-1-172.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
1-172.pdf.

Anonymous:1936:ATAb

[Ano36b] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 15(2):349–350, April 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-2-349.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
2-349.pdf.

Anonymous:1936:ATAc

[Ano36c] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 15(3):476–479, July 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-3-476.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
3-476.pdf.

Anonymous:1936:ATAd

[Ano36d] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 15(4):628–632, October 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-4-628.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
4-628.pdf.

Anonymous:1936:CIa

[Ano36e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 15(1):177–179, January 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-1-177.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
1-177.pdf.

Anonymous:1936:CIb

[Ano36f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 15(2):351–353, April 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-2-351.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        2-351.pdf.

Anonymous:1936:CIc

[Ano36g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 15(3):480–482, July 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-3-480.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        3-480.pdf.

Anonymous:1936:CId

[Ano36h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 15(4):633–634, October 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-4-633.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        4-633.pdf.

Anonymous:1936:TDU

[Ano36i] Anonymous. Technical developments underlying the toll services of the Bell System: A supplement to the Bell System Technical Journal, July 1936. The Bell System Technical Journal, 15(4):1–80, October 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-4-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol15-
1936/articles/bstj15-4-1.pdf.

Anonymous:1937:ATAa

[Ano37a] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 16(1): 113–115, January 1937. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-1-113.pdf;   http://
www.alcatel-lucent.com/bstj/
vol16-1937/articles/bstj16-
1-113.pdf.

Anonymous:1937:ATAb

[Ano37b] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 16(4):578–580, October 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-4-578.pdf;   http://
www.alcatel-lucent.com/bstj/
vol16-1937/articles/bstj16-
4-578.pdf.

Anonymous:1937:ATPa

[Ano37c] Anonymous. Abstracts of technical papers from Bell System sources. The Bell System Technical Journal, 16(2):247–248, April 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-2-247.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        2-247.pdf.

Anonymous:1937:ATPb

[Ano37d] Anonymous. Abstracts of technical papers from Bell System sources. The Bell System Technical Journal, 16(3):420–421, July 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-3-420.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        3-420.pdf.

Anonymous:1937:CIc

[Ano37e] Anonymous. Contributors in this issue. The Bell System Technical Journal, 16(3):422, July 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-3-422.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        3-422.pdf.

Anonymous:1937:CIa

[Ano37f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 16(1):116–117, January 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-1-116.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol16-1937/articles/bstj16-
1-116.pdf.

Anonymous:1937:CIb

[Ano37g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 16(2):249–250, April 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-2-249.pdf;   http://
www.alcatel-lucent.com/bstj/
vol16-1937/articles/bstj16-
2-249.pdf.

Anonymous:1937:CId

[Ano37h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 16(4):581–583, October 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-4-581.pdf;   http://
www.alcatel-lucent.com/bstj/
vol16-1937/articles/bstj16-
4-581.pdf.

Anonymous:1938:ATAa

[Ano38a] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 17(1): 184–190, January 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-1-184.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
1-184.pdf.

Anonymous:1938:ATAb

[Ano38b] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 17(2):319–323, April 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-2-319.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        2-319.pdf.

Anonymous:1938:ATAc

[Ano38c] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 17(3):483–485, July 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-3-483.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        3-483.pdf.

Anonymous:1938:ATAd

[Ano38d] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 17(4):670–672, October 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-4-670.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        4-670.pdf.

Anonymous:1938:CIa

[Ano38e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 17(1):191–193, January 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-1-191.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
1-191.pdf.

Anonymous:1938:CIb

[Ano38f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 17(2):324–325, April 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-2-324.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
2-324.pdf.

Anonymous:1938:CIc

[Ano38g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 17(3):486–488, July 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-3-486.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
3-486.pdf.

Anonymous:1938:CId

[Ano38h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 17(4):673–674, October 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-4-673.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
4-673.pdf.

Anonymous:1939:ATAa

[Ano39a] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 18(1): 246–250, January 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-1-246.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
1-246.pdf.

Anonymous:1939:ATAb

[Ano39b] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 18(2):388–390, April 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-2-388.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
2-388.pdf.

Anonymous:1939:ATAc

[Ano39c] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 18(3):538–543, July 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-3-538.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
3-538.pdf.

Anonymous:1939:ATAd

[Ano39d] Anonymous. Abstracts of technical articles from Bell System sources. The Bell System Technical Journal, 18(4):742–746, October 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-4-742.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
4-742.pdf.

Anonymous:1939:CIa

[Ano39e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 18(1):251–254, January 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-1-251.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
1-251.pdf.

Anonymous:1939:CIb

[Ano39f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 18(2):391–393, April 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/
                                        bstj18-2-391.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol18-1939/articles/bstj18-
                                        2-391.pdf.

Anonymous:1939:CIc

[Ano39g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 18(3):544–545, July 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/
                                        bstj18-3-544.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol18-1939/articles/bstj18-
                                        3-544.pdf.

Anonymous:1939:CId

[Ano39h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 18(4):747–748, October 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/
                                        bstj18-4-747.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol18-1939/articles/bstj18-
                                        4-747.pdf.

Blattner:1930:STS

[BB30] D. G. Blattner and L. G. Bostwick. Sound transmission system for two-way television. The Bell System Technical Journal, 9(3):478–482, July 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-

3-478.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-3-478.pdf.

Burton:1933:AFA

[BB33] E. T. Burton and E. M. Boardman. Audio frequency atmospherics. The Bell System Technical Journal, 12(4):498–516, October 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-4-498.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
4-498.pdf.

Bruce:1935:EDS

[BB35] E. Bruce and A. C. Beck. Experiments with directivity steering for fading reduction. The Bell System Technical Journal, 14 (2):195–210, April 1935. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-2-195.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
2-195.pdf.

Barstow:1936:TDM

[BBK36] J. M. Barstow, P. W. Blye, and H. E. Kent. Technical digests– measurement of telephone noise and power wave shape. The Bell System Technical Journal, 15(1): 151–156, January 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-1-151.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        1-151.pdf.

Bruce:1935:HRA

[BBL35] E. Bruce, A. C. Beck, and L. R. Lowry. Horizontal rhombic antennas. The Bell System Technical Journal, 14(1):135–158, January 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-1-135.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        1-135.pdf.

Bode:1935:IWF

[BD35] H. W. Bode and R. L. Dietzold. Ideal wave filters. The Bell System Technical Journal, 14 (2):215–252, April 1935. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-2-215.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        2-215.pdf.

Becker:1935:TEE

[Bec35] J. A. Becker. Thermionic electron emission. The Bell System Technical Journal, 14(3):413–476, July 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-3-413.pdf;   http://

www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
3-413.pdf.

Bennett:1933:NRC

[Ben33] W. R. Bennett. New results in the calculation of modulation products. The Bell System Technical Journal, 12(2):228–243, April 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-2-228.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
2-228.pdf.

Best:1933:RTM

[Bes33] F. H. Best. A recording transmission measuring system for telephone circuit testing. The Bell System Technical Journal, 12 (1):22–34, January 1933. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-1-22.pdf;   http://www.
alcatel-lucent.com/bstj/vol12-
1933/articles/bstj12-1-22.pdf.

Best:1938:NTM

[Bes38] F. H. Best. New transmission measuring systems for telephone circuit maintenance. The Bell System Technical Journal, 17(1):1–16, January 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-1-1.pdf;   http://www.

                                        alcatel-lucent.com/bstj/vol17-
                                        1938/articles/bstj17-1-1.pdf.

Bowen:1930:MIG

[BG30] A. E. Bowen and G. L. Gilkeson. Mutual impedances of ground-return circuits. The Bell System Technical Journal, 9(4): 628–651, October 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        4-628.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-4-628.pdf.

Bailey:1936:LWR

[BH36a] Austin Bailey and A. E. Harper. Long-wave radio transmission phenomena associated with a cessation of the Sun’s rays. The Bell System Technical Journal, 15(1):1–19, January 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-1-1.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol15-
                                        1936/articles/bstj15-1-1.pdf.

Burns:1936:TDD

[BH36b] R. M. Burns and H. E. Haring. Technical digests—determination of the corrosion behavior of painted iron and the inhibitive action of paints. The Bell System Technical Journal, 15 (2):343–348, April 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol15/
bstj15-2-343.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
2-343.pdf.

Burrows:1935:USW

[BHD35] C. R. Burrows, L. E. Hunt, and A. Decino. Ultra-short-wave propagation: Mobile urban transmission characteristics. The Bell System Technical Journal, 14(2):253–272, April 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-2-253.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
2-253.pdf.

Bedell:1934:SAS

[BK34] E. H. Bedell and Iden Kerney. System adaptation, symposium on wire transmission of symphonic music and its reproduction in auditory perspective. The Bell System Technical Journal, 13(2):301–308, April 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-2-301.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
2-301.pdf.

Blackwell:1932:TFT

[Bla32] O. B. Blackwell. The time factor in telephone transmission. The Bell System Technical Journal, 11(1):53–66, January 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol11/
                                        bstj11-1-53.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol11-
                                        1932/articles/bstj11-1-53.pdf.

Black:1934:SFA

[Bla34] H. S. Black. Stabilized feedback amplifiers. The Bell System Technical Journal, 13(1):1–18, January 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol13/
                                        bstj13-1-1.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol13-
                                        1934/articles/bstj13-1-1.pdf.

Blanchard:1938:HDE

[Bla38] Julian Blanchard. Hertz, the discoverer of electric waves. The Bell System Technical Journal, 17(3):327–337, July 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-3-327.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        3-327.pdf.

Bailey:1931:APT

[BM31] Austin Bailey and T. A. Mc-Cann. Application of printing telegraph to long-wave radio circuits. The Bell System Technical Journal, 10(4):601–615, October 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-4-601.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
4-601.pdf.

Bode:1930:MIC

[Bod30] H. W. Bode. A method of impedance correction. The Bell System Technical Journal, 9(4): 794–835, October 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
4-794.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-4-794.pdf.

Bode:1935:GTE

[Bod35] H. W. Bode. A general theory of electric wave filters. The Bell System Technical Journal, 14 (2):211–214, April 1935. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-2-211.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
2-211.pdf.

Bode:1938:VE

[Bod38] H. W. Bode. Variable equalizers. The Bell System Technical Journal, 17(2):229–244, April 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-2-229.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        2-229.pdf.

Bown:1930:TTS

[Bow30] Ralph Bown. Transoceanic telephone service — short-wave transmission. The Bell System Technical Journal, 9(2): 258–269, April 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        2-258.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-2-258.pdf.

Bown:1937:TDT

[Bow37] Ralph Bown. Technical digests — transoceanic radio telephone development. The Bell System Technical Journal, 16(4): 560–567, October 1937. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-4-560.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        4-560.pdf.

Bozorth:1936:PSF

[Boz36] R. M. Bozorth. The present status of ferromagnetic theory. The Bell System Technical Journal, 15(1):63–91, January 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-1-63.pdf;   http://www.

alcatel-lucent.com/bstj/vol15-
1936/articles/bstj15-1-63.pdf.

Bronson:1936:TOB

[Bro36] F. M. Bronson. Tandem operation in the Bell System. The Bell System Technical Journal, 15(3):380–404, July 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-3-380.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
3-380.pdf.

Bedell:1939:EAC

[BRS39] W. B. Bedell, G. B. Ransom, and W. A. Stevens. Experience in applying carrier telephone systems to toll cables. The Bell System Technical Journal, 18(4): 547–587, October 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-4-547.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
4-547.pdf.

Bruce:1931:DSW

[Bru31] E. Bruce. Developments in shortwave directive antennas. The Bell System Technical Journal, 10 (4):656–683, October 1931. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-4-656.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        4-656.pdf.

Bailey:1935:TLW

[BT35] Austin Bailey and Howard M. Thomson. Transatlantic long-wave radio telephone transmission and related phenomena from 1923 to 1933. The Bell System Technical Journal, 14(4):680–697, October 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-4-680.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        4-680.pdf.

Burrows:1935:RPS

[Bur35] Charles R. Burrows. Radio propagation over spherical Earth. The Bell System Technical Journal, 14(3):477–488, July 1935. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-3-477.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        3-477.pdf.

Burns:1936:CMM

[Bur36a] R. M. Burns. The corrosion of metals—I. mechanism of corrosion processes. The Bell System Technical Journal, 15(1): 20–38, January 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol15/
bstj15-1-20.pdf;   http://www.
alcatel-lucent.com/bstj/vol15-
1936/articles/bstj15-1-20.pdf.

Burns:1936:CMI

[Bur36b] R. M. Burns. Corrosion of metals—II. lead and lead-alloy cable sheathing. The Bell System Technical Journal, 15(4): 603–625, October 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-4-603.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
4-603.pdf.

Burrows:1937:AQR

[Bur37a] Charles R. Burrows. Addendum to ‘Radio Propagation Over Plane Earth — Field Strength Curves’. The Bell System Technical Journal, 16(4):574–577, October 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-4-574.pdf;   http://
www.alcatel-lucent.com/bstj/
vol16-1937/articles/bstj16-
4-574.pdf.

Burrows:1937:RPP

[Bur37b] Charles R. Burrows. Radio propagation over plane Earth — field strength curves. The Bell System Technical Journal, 16 (1):45–75, January 1937. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-1-45.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol16-
                                        1937/articles/bstj16-1-45.pdf.

Burrows:1938:ETL

[Bur38] Charles R. Burrows. The exponential transmission line. The Bell System Technical Journal, 17 (4):555–573, October 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-4-555.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        4-555.pdf.

Colley:1936:RBP

[CA36] R. H. Colley and C. H. Amadon. The relation between penetration and decay in creosoted southern pine poles. The Bell System Technical Journal, 15(3):363–379, July 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-3-363.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        3-363.pdf.

Carson:1930:NHO

[Car30a] John R. Carson. Notes on the Heaviside operational calculus. The Bell System Technical Journal, 9(1):150–162, January 1930. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol09/
bstj9-1-150.pdf;   http://www.
alcatel-lucent.com/bstj/vol09-
1930/articles/bstj9-1-150.pdf.

Carson:1930:RET

[Car30b] John R. Carson. The reciprocal energy theorem. The Bell System Technical Journal, 9 (2):325–331, April 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
2-325.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-2-325.pdf.

Carson:1931:SEF

[Car31] John R. Carson. The statistical energy-frequency spectrum of random disturbances. The Bell System Technical Journal, 10(3):374–381, July 1931. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-3-374.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
3-374.pdf.

Card:1936:TDE

[Car36a] R. H. Card. Technical digests— Earth resistivity and geological structure. The Bell System Technical Journal, 15(1): 167–171, January 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/

                                        bstj15-1-167.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        1-167.pdf.

Carr:1936:FOW

[Car36b] J. A. Carr. Forces of oblique winds on telephone wires. The Bell System Technical Journal, 15 (4):587–602, October 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-4-587.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        4-587.pdf.

Carson:1936:EOC

[Car36c] John R. Carson. An extension of operational calculus. The Bell System Technical Journal, 15 (2):340–342, April 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-2-340.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        2-340.pdf.

Carson:1939:FMT

[Car39a] John R. Carson. Frequencymodulation: Theory of the feedback receiving circuit. The Bell System Technical Journal, 18 (3):395–403, July 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/

bstj18-3-395.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
3-395.pdf.

Caruthers:1939:COM

[Car39b] R. S. Caruthers. Copper oxide modulators in carrier telephone systems. The Bell System Technical Journal, 18(2):315–337, April 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-2-315.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
2-315.pdf.

Castner:1933:DAA

[CC33] T. G. Castner and C. W. Carter, Jr. Developments in the application of articulation testing. The Bell System Technical Journal, 12(3):347–370, July 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-3-347.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
3-347.pdf.

Coleman:1937:ICC

[CD37] J. O’R. Coleman and R. F. Davis.

The inductive coordination of common-neutral power distribution systems and telephone circuits. The Bell System Technical Journal, 16(1):76–100, January 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-1-76.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol16-
                                        1937/articles/bstj16-1-76.pdf.

Carson:1937:VFE

[CF37] John R. Carson and Thornton C. Fry. Variable frequency electric circuit theory with application to the theory of frequencymodulation. The Bell System Technical Journal, 16(4): 513–540, October 1937. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-4-513.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        4-513.pdf.

Clark:1930:LDC

[CG30] A. B. Clark and C. W. Green. Long distance cable circuit for program transmission. The Bell System Technical Journal, 9(3):567–594, July 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        3-567.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-3-567.pdf.

Clarke:1936:MSM

[CH36] Beverly L. Clarke and H. W. Hermance. Microchemical and special methods of analysis in communication research. The Bell System Technical Journal, 15(4):

483–503, October 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-4-483.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
4-483.pdf.

Carr:1938:STL

[CH38] J. A. Carr and F. V. Haskell.

Studies of telephone line wire spacing problems. The Bell System Technical Journal, 17 (2):195–228, April 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-2-195.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
2-195.pdf.

Chapman:1934:OWCa

[Cha34a] A. G. Chapman. Open-wire crosstalk. The Bell System Technical Journal, 13(1):19–58, January 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-1-19.pdf;   http://www.
alcatel-lucent.com/bstj/vol13-
1934/articles/bstj13-1-19.pdf.

Chapman:1934:OWCb

[Cha34b] A. G. Chapman. Open-wire crosstalk. The Bell System Technical Journal, 13(2):195–238, April 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol13/
                                        bstj13-2-195.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol13-1934/articles/bstj13-
                                        2-195.pdf.

Chaffee:1939:ANF

[Cha39] J. G. Chaffee. The application of negative feedback to frequencymodulation systems. The Bell System Technical Journal, 18 (3):404–437, July 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/
                                        bstj18-3-404.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol18-1939/articles/bstj18-
                                        3-404.pdf.

Christopher:1932:TCC

[Chr32] A. J. Christopher. Transformer coupling circuits for highfrequency amplifiers. The Bell System Technical Journal, 11(4): 608–621, October 1932. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol11/
                                        bstj11-4-608.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol11-1932/articles/bstj11-
                                        4-608.pdf.

Chesnut:1938:CCT

[CIK38] R. W. Chesnut, L. M. Ilgenfritz, and A. Kenner. Cable carrier telephone terminals. The Bell System Technical Journal, 17(1): 106–124, January 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-1-106.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
1-106.pdf.

Clark:1933:CC

[CK33] A. B. Clark and B. W. Kendall.

Carrier in cable. The Bell System Technical Journal, 12 (3):251–263, July 1933. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-3-251.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
3-251.pdf.

Clark:1930:SRD

[Cla30a] A. B. Clark. Some recent developments in long distance cables in the United States of America. The Bell System Technical Journal, 9(3):487–492, July 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
3-487.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-3-487.pdf.

Clark:1930:WLS

[Cla30b] A. B. Clark. Wire line systems for national broadcasting. The Bell System Technical Journal, 9 (1):141–149, January 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        1-141.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-1-141.pdf.

Clark:1935:WBT

[Cla35] A. B. Clark. Wide band transmission over balanced circuits. The Bell System Technical Journal, 14(1):1–7, January 1935. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-1-1.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol14-
                                        1935/articles/bstj14-1-1.pdf.

Clement:1934:LFP

[Cle34]  A. W. Clement.  Line filter

for program system. The Bell System Technical Journal, 13 (3):382–390, July 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol13/
                                        bstj13-3-382.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol13-1934/articles/bstj13-
                                        3-382.pdf.

Carson:1936:HFW

[CMS36] John R. Carson, Sallie P. Mead, and S. A. Schelkunoff. Hyper-frequency wave guides— mathematical theory. The Bell System Technical Journal, 15 (2):310–333, April 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-2-310.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
2-310.pdf.

Clark:1932:LDT

[CO32] A. B. Clark and H. S. Osborne.

Long distance telephone circuits in cable. The Bell System Technical Journal, 11(4):520–545, October 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-4-520.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
4-520.pdf.

Colpitts:1937:RTT

[Col37a] Edwin H. Colpitts. Recent trends in toll transmission in the United States. The Bell System Technical Journal, 16(2):119–143, April 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-2-119.pdf;   http://
www.alcatel-lucent.com/bstj/
vol16-1937/articles/bstj16-
2-119.pdf.

Colpitts:1937:SRA

[Col37b] Edwin H. Colpitts. Scientific research applied to the telephone transmitter and receiver. The Bell System Technical Journal, 16(3):251–274, July 1937. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-3-251.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        3-251.pdf.

Christensen:1936:SRF

[CP36] C. J. Christensen and G. L. Pearson. Spontaneous resistance fluctuations in carbon microphones and other granular resistances. The Bell System Technical Journal, 15(2):197–223, April 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-2-197.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        2-197.pdf.

Clarke:1938:SAC

[CR38] Beverly L. Clarke and A. E. Ruehle. Spectrochemical analysis in communication research. The Bell System Technical Journal, 17(3):381–392, July 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-3-381.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        3-381.pdf.

Crisson:1931:NIT

[Cri31]  George Crisson.  Negative

impedances and the twin 21-type repeater. The Bell System Technical Journal, 10(3):485–513, July 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-3-485.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
3-485.pdf.

Curtis:1933:OTT

[Cur33] A. M. Curtis. An oscillograph for ten thousand cycles. The Bell System Technical Journal, 12 (1):76–90, January 1933. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-1-76.pdf;   http://www.
alcatel-lucent.com/bstj/vol12-
1933/articles/bstj12-1-76.pdf.

Conwell:1931:SJD

[CW31] R. N. Conwell and H. S. Warren.

Status of joint development and research on low-frequency induction. The Bell System Technical Journal, 10(2):206–230, April 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-2-206.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
2-206.pdf.

Darrow:1930:CAPa

[Dar30a] Karl K. Darrow. Contemporary advances in physics, XIX, fusion of wave and corpuscle theories. The Bell System Technical Journal, 9 (1):163–188, January 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        1-163.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-1-163.pdf.

Darrow:1930:CAPb

[Dar30b] Karl K. Darrow. Contemporary advances in physics, XX, ionization of gases by light. The Bell System Technical Journal, 9 (2):341–355, April 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        2-341.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-2-341.pdf.

Darrow:1930:CAPc

[Dar30c] Karl K. Darrow. Contemporary advances in physics, XXI, interception and scattering of electrons and ions. The Bell System Technical Journal, 9(4): 668–696, October 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        4-668.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-4-668.pdf.

Darrow:1931:CAP

[Dar31] Karl K. Darrow. Contemporary advances in physics, XXII — transmutation. The Bell System Technical Journal, 10(4): 628–655, October 1931. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-4-628.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
4-628.pdf.

Darrow:1932:CAPa

[Dar32a] Karl K. Darrow. Contemporary advances in physics, XXIII — data and nature of cosmic rays. The Bell System Technical Journal, 11(1):148–184, January 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-1-148.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
1-148.pdf.

Darrow:1932:CAPb

[Dar32b] Karl K. Darrow. Contemporary advances in physics, XXIV — high-frequency phenomena in gases, first part. The Bell System Technical Journal, 11(4): 576–607, October 1932. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-4-576.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
4-576.pdf.

Darrow:1933:CAPa

[Dar33a] Karl K. Darrow. Contemporary advances in physics, XXV, highfrequency phenomena in gases, second part. The Bell System Technical Journal, 12(1): 91–118, January 1933. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-1-91.pdf;   http://www.
alcatel-lucent.com/bstj/vol12-
1933/articles/bstj12-1-91.pdf.

Darrow:1933:CAPb

[Dar33b] Karl K. Darrow. Contemporary advances in physics, XXVI, the nucleus, first part. The Bell System Technical Journal, 12(3):288–330, July 1933. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-3-288.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
3-288.pdf.

Darrow:1934:CAPa

[Dar34a] Karl K. Darrow. Contemporary advances in physics, XXVII the nucleus, second part. The Bell System Technical Journal, 13(1): 102–158, January 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-1-102.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
1-102.pdf.

Darrow:1934:CAPb

[Dar34b] Karl K. Darrow. Contemporary advances in physics, XXVIII — the nucleus, third part. The Bell System Technical Journal, 13(3):391–404, July 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-3-391.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
3-391.pdf.

Darrow:1934:CAPc

[Dar34c] Karl K. Darrow. Contemporary advances in physics, XXVIII — the nucleus, third part. The Bell System Technical Journal, 13 (4):580–613, October 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-4-580.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
4-580.pdf.

Darrow:1935:CAP

[Dar35] Karl K. Darrow. Contemporary advances in physics, XXIX, the nucleus, fourth part. The Bell System Technical Journal, 14 (2):285–321, April 1935. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-2-285.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol14-1935/articles/bstj14-
                                        2-285.pdf.

Darrow:1936:CAP

[Dar36] Karl K. Darrow. Contemporary advances in physics, XXX, the theory of magnetism. The Bell System Technical Journal, 15 (2):224–247, April 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-2-224.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        2-224.pdf.

Darrow:1937:CAP

[Dar37] Karl K. Darrow. Contemporary advances in physics, XXXI — spinning atoms and spinning electrons. The Bell System Technical Journal, 16(3):319–336, July 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-3-319.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        3-319.pdf.

Darrow:1938:RAN

[Dar38] Karl K. Darrow. Radioactivity — artificial and natural. The Bell System Technical Journal, 17 (2):292–318, April 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-2-292.pdf;   http://

www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
2-292.pdf.

Darrow:1939:CAP

[Dar39] Karl K. Darrow. Contemporary advances in physics, XXXII, particles of the cosmic rays. The Bell System Technical Journal, 18 (1):190–217, January 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-1-190.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
1-190.pdf.

Davisson:1932:CDE

[Dav32] C. J. Davisson. The conception and demonstration of electron waves. The Bell System Technical Journal, 11(4):546–562, October 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-4-546.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
4-546.pdf.

Davisson:1938:DEW

[Dav38] C. J. Davisson. The discovery of electron waves. The Bell System Technical Journal, 17(3):475–482, July 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-3-475.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        3-475.pdf.

Dietze:1939:CCN

[DG39] E. Dietze and W. D. Goodale, Jr. The computation of the composite noise resulting from random variable sources. The Bell System Technical Journal, 18(4): 605–623, October 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/
                                        bstj18-4-605.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol18-1939/articles/bstj18-
                                        4-605.pdf.

Davis:1933:STP

[DH33] R. F. Davis and H. R. Huntley. Some theoretical and practical aspects of noise induction. The Bell System Technical Journal, 12(4):469–497, October 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol12/
                                        bstj12-4-469.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol12-1933/articles/bstj12-
                                        4-469.pdf.

Doherty:1936:TDN

[Doh36] W. H. Doherty. Technical digests—A new high-efficiency power amplifier for modulated waves. The Bell System Technical Journal, 15(3):469–475, July 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-3-469.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
3-469.pdf.

Englund:1933:SRS

[ECM33] C. R. Englund, A. B. Crawford, and W. W. Mumford. Some results of a study of ultra-short-wave transmission phenomena. The Bell System Technical Journal, 12(2):197–227, April 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-2-197.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
2-197.pdf.

Englund:1935:FRS

[ECM35] C. R. Englund, A. B. Crawford, and W. W. Mumford. Further results of a study of ultrashort-wave transmission phenomena. The Bell System Technical Journal, 14(3):369–387, July 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-3-369.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
3-369.pdf.

Englund:1938:USW

[ECM38] C. R. Englund, A. B. Crawford, and W. W. Mumford. Ultrashort-wave transmission and atmospheric irregularities. The Bell System Technical Journal, 17(4): 489–519, October 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-4-489.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        4-489.pdf.

Ellwood:1937:SML

[EL37] W. B. Ellwood and V. E. Legg. Study of magnetic losses at low flux densities in permalloy sheet. The Bell System Technical Journal, 16(2):212–227, April 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-2-212.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        2-212.pdf.

Elmen:1936:MAI

[Elm36] G. W. Elmen. Magnetic alloys of iron, nickel, and cobalt. The Bell System Technical Journal, 15(1):113–135, January 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-1-113.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        1-113.pdf.

Espenschied:1939:TCI

[EN39] Lloyd Espenschied and R. C. Newhouse. A terrain clearance indicator. The Bell System Technical Journal, 18(1):222–234, January 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-1-222.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
1-222.pdf.

Espenschied:1934:SWB

[ES34] L. Espenschied and M. E. Strieby.

Systems for wide-band transmission over coaxial lines. The Bell System Technical Journal, 13(4): 654–679, October 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-4-654.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
4-654.pdf.

Ellis:1935:SMM

[ES35] W. C. Ellis and Earle E. Schumacher. A survey of magnetic materials in relation to structure. The Bell System Technical Journal, 14(1):8–43, January 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-1-8.pdf;   http://www.
alcatel-lucent.com/bstj/vol14-
1935/articles/bstj14-1-8.pdf.

Espenschied:1931:ORE

[EW31] Lloyd Espenschied and William Wilson. Overseas radio extensions to wire telephone networks. The Bell System Technical Journal, 10(2):243–264, April 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-2-243.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
2-243.pdf.

Fisher:1938:NSC

[FAM38] H. J. Fisher, M. L. Almquist, and R. H. Mills. A new single channel carrier telephone system. The Bell System Technical Journal, 17(1):162–183, January 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-1-162.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
1-162.pdf.

Fay:1932:OVT

[Fay32] C. E. Fay. The operation of vacuum tubes as class B and class C amplifiers. The Bell System Technical Journal, 11 (1):28–52, January 1932. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-1-28.pdf;   http://www.
alcatel-lucent.com/bstj/vol11-
1932/articles/bstj11-1-28.pdf.

French:1930:WST

[FCK30] Norman R. French, Charles W.

Carter, Jr., and Walter Koenig, Jr. The words and sounds of telephone conversations. The Bell System Technical Journal, 9 (2):290–324, April 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
2-290.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-2-290.pdf.

Ferguson:1933:CBM

[Fer33] John G. Ferguson. Classification of bridge methods of measuring impedances. The Bell System Technical Journal, 12(4): 452–468, October 1933. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-4-452.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
4-452.pdf.

Friis:1937:MUS

[FF37] H. T. Friis and C. B. Feldman. A multiple unit steerable antenna for short-wave reception. The Bell System Technical Journal, 16(3):337–419, July 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-3-337.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol16-1937/articles/bstj16-
                                        3-337.pdf.

Ferris:1936:TDE

[FKSW36] L. P. Ferris, B. G. King, P. W. Spence, and H. B. Williams. Technical digests—effect of electric shock on the heart. The Bell System Technical Journal, 15(3):455–468, July 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-3-455.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        3-455.pdf.

Flanders:1932:MMA

[Fla32] P. B. Flanders. A method of measuring acoustic impedance. The Bell System Technical Journal, 11(3):402–410, July 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol11/
                                        bstj11-3-402.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol11-1932/articles/bstj11-
                                        3-402.pdf.

Fletcher:1931:SPC

[Fle31] Harvey Fletcher. Some physical characteristics of speech and music. The Bell System Technical Journal, 10(3):349–373, July 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-3-349.pdf;   http://

www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
3-349.pdf.

Fletcher:1934:SWT

[Fle34] Harvey Fletcher. Symposium on wire transmission of symphonic music and its reproduction in auditory perspective, basic requirements. The Bell System Technical Journal, 13(2):239–244, April 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-2-239.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
2-239.pdf.

Fletcher:1933:LDM

[FM33] Harvey Fletcher and W. A. Munson. Loudness, its definition, measurement and calculation. The Bell System Technical Journal, 12 (4):377–430, October 1933. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-4-377.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
4-377.pdf.

Fondiller:1930:DCM

[Fon30] William Fondiller. Developments in communication materials. The Bell System Technical Journal, 9 (2):237–257, April 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.

                                        com/BSTJ/images/Vol09/bstj9-
                                        2-237.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-2-237.pdf.

Foster:1931:MIG

[Fos31]  Ronald M. Foster.  Mutual

impedance of grounded wires lying on the surface of the Earth. The Bell System Technical Journal, 10(3):408–419, July 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-3-408.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        3-408.pdf.

Foster:1933:MIG

[Fos33]  Ronald M. Foster.  Mutual

impedance of grounded wires lying on or above the surface of the Earth. The Bell System Technical Journal, 12(3):264–287, July 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol12/
                                        bstj12-3-264.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol12-1933/articles/bstj12-
                                        3-264.pdf.

Fruth:1932:CSC

[Fru32] Hal F. Fruth. Cathode sputtering — A commercial application. The Bell System Technical Journal, 11(2):283–292, April 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol11/
bstj11-2-283.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
2-283.pdf.

Fay:1938:TSC

[FSS38] C. E. Fay, A. L. Samuel, and W. Shockley. On the theory of space charge between parallel plane electrodes. The Bell System Technical Journal, 17 (1):49–79, January 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-1-49.pdf;   http://www.
alcatel-lucent.com/bstj/vol17-
1938/articles/bstj17-1-49.pdf.

Green:1938:CTS

[GG38] C. W. Green and E. I. Green. A carrier telephone system for toll cables. The Bell System Technical Journal, 17(1):80–105, January 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-1-80.pdf;   http://www.
alcatel-lucent.com/bstj/vol17-
1938/articles/bstj17-1-80.pdf.

Gherardi:1931:SCP

[Ghe31] B. Gherardi. Symposium on coordination of power and telephone plant, closing remarks. The Bell System Technical Journal, 10(2):241–242, April 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-2-241.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        2-241.pdf.

Gherardi:1932:HEP

[Ghe32] Bancroft Gherardi. Henry as an electrical pioneer. The Bell System Technical Journal, 11 (3):327–333, July 1932. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol11/
                                        bstj11-3-327.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol11-1932/articles/bstj11-
                                        3-327.pdf.

Gibbon:1938:ECB

[Gib38] C. O. Gibbon. An explanation of the common battery antisidetone subscriber set. The Bell System Technical Journal, 17 (2):245–257, April 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-2-245.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        2-245.pdf.

Gillett:1931:SDC

[Gil31] G. D. Gillett. Some developments in common frequency broadcasting. The Bell System Technical Journal, 10(4):577–600, October 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol10/
bstj10-4-577.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
4-577.pdf.

Gherardi:1930:TCS

[GJ30] Bancroft Gherardi and F. B. Jewett. Telephone communication system of the United States. The Bell System Technical Journal, 9 (1):1–100, January 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol09/
bstj9-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol09-
1930/articles/bstj9-1-1.pdf.

Gherardi:1932:WWT

[GJ32] Bancroft Gherardi and Frank B.

Jewett. World-wide telephony — its problems and future. The Bell System Technical Journal, 11 (4):485–519, October 1932. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-4-485.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
4-485.pdf.

Ganz:1936:ICT

[GL36] A. G. Ganz and A. G. Laird.

Improvements in communication transformers. The Bell System Technical Journal, 15(1): 136–150, January 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-1-136.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        1-136.pdf.

Green:1936:PSC

[GLC36] E. I. Green, F. A. Leibe, and H. E. Curtis. The proportioning of shielded circuits for minimum high-frequency attenuation. The Bell System Technical Journal, 15(2):248–283, April 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-2-248.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        2-248.pdf.

Gifford:1935:TDM

[GM35] F. A. Gifford and R. B. Meader. Technical digests: Marine radio telephone service for Boston Harbor. The Bell System Technical Journal, 14(4):702–707, October 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-4-702.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        4-702.pdf.

Goucher:1934:CMA

[Gou34] F. S. Goucher. The carbon microphone: An account of some researches bearing on its action. The Bell System Technical Journal, 13(2):163–194, April 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-2-163.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
2-163.pdf.

Gray:1939:EEO

[Gra39] Frank Gray. Electrostatic electronoptics. The Bell System Technical Journal, 18(1):1–31, January 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol18-
1939/articles/bstj18-1-1.pdf.

Green:1930:TCO

[Gre30] E. I. Green. The transmission characteristics of open-wire telephone lines. The Bell System Technical Journal, 9(4): 730–759, October 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
4-730.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-4-730.pdf.

Gustafson:1938:MST

[Gus38] W. G. Gustafson. Magnetic shielding of transformers at audio frequencies. The Bell System Technical Journal, 17(3):416–437, July 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-3-416.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        3-416.pdf.

Hamilton:1934:WBO

[Ham34] H. S. Hamilton. Wide-band open-wire program system. The Bell System Technical Journal, 13(3):351–381, July 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol13/
                                        bstj13-3-351.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol13-1934/articles/bstj13-
                                        3-351.pdf.

Hansen:1935:TDH

[Han35] E. B. Hansen. Technical digests: Harbor craft ship-to-shore radio telephone service in puget sound area. The Bell System Technical Journal, 14(4):708–712, October 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-4-708.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        4-708.pdf.

Harris:1932:PMU

[Har32] G. R. Harris. Precision methods used in constructing electric wave filters for carrier systems. The Bell System Technical Journal, 11(2):264–282, April 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-2-264.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
2-264.pdf.

Hartley:1936:OSN

[Har36] R. V. L. Hartley. Oscillations in systems with non-linear reactance. The Bell System Technical Journal, 15(3):424–440, July 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-3-424.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
3-424.pdf.

Harrison:1939:HFS

[Har39] W. H. Harrison. Hurricane and flood — September 1938. The Bell System Technical Journal, 18 (1):218–221, January 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-1-218.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
1-218.pdf.

Haworth:1931:MCT

[Haw31] F. E. Haworth. A magnetic curve tracer. The Bell System Technical Journal, 10(1):20–32, January 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-1-20.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol10-
                                        1931/articles/bstj10-1-20.pdf.

Hunter:1935:CCE

[HB35] R. N. Hunter and R. P. Booth. Cable crosstalk–effect of non-uniform current distribution in the wires. The Bell System Technical Journal, 14(2):179–194, April 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-2-179.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        2-179.pdf.

Holbrook:1939:LRT

[HD39] B. D. Holbrook and J. T. Dixon. Load rating theory for multichannel amplifiers. The Bell System Technical Journal, 18(4): 624–644, October 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/
                                        bstj18-4-624.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol18-1939/articles/bstj18-
                                        4-624.pdf.

Henneberger:1931:BML

[HE31] T. C. Henneberger and P. G. Edwards. Bridge methods for locating resistance faults on cable wires. The Bell System Technical Journal, 10(3):382–407, July 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-3-382.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
3-382.pdf.

Herriott:1938:HSM

[Her38] W. Herriott. High speed motion picture photography. The Bell System Technical Journal, 17(3):393–405, July 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-3-393.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
3-393.pdf.

Harrison:1932:EMC

[HF32] H. C. Harrison and P. B. Flanders. An efficient miniature condenser microphone system. The Bell System Technical Journal, 11(3):451–461, July 1932. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-3-451.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
3-451.pdf.

Hickman:1937:SRM

[Hic37] C. N. Hickman. Sound recording on magnetic tape. The Bell System Technical Journal, 16 (2):165–177, April 1937. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-2-165.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        2-165.pdf.

Hoyt:1935:MIP

[HM35] Ray S. Hoyt and Sallie Pero Mead. Mutual impedances of parallel wires. The Bell System Technical Journal, 14(3):509–533, July 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-3-509.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        3-509.pdf.

Huntley:1935:SAL

[HO35] H. R. Huntley and E. J. O’Connell. Some aspects of low-frequency induction between power and telephone circuits. The Bell System Technical Journal, 14(4): 573–599, October 1935. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-4-573.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        4-573.pdf.

Horton:1938:OEL

[Hor38] Arthur W. Horton, Jr. The occurrence and effect of lockout occasioned by two echo suppressors. The Bell System Technical Journal, 17(2):258–280, April 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-2-258.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
2-258.pdf.

Hoyt:1933:PTT

[Hoy33] Ray S. Hoyt. Probability theory and telephone transmission engineering. The Bell System Technical Journal, 12(1):35–75, January 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-1-35.pdf;   http://www.
alcatel-lucent.com/bstj/vol12-
1933/articles/bstj12-1-35.pdf.

Harrison:1931:TTP

[HS31] W. H. Harrison and A. E. Silver. Trends in telephone and power practise as affecting coordination. The Bell System Technical Journal, 10(2):159–183, April 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-2-159.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
2-159.pdf.

Haring:1935:TDS

[HTSP35] H. E. Haring, U. B. Thomas, E. E. Schumacher, and G. S. Phipps. Technical digests: Superiorities of lead-calcium alloys for storage battery construction.

The Bell System Technical Journal, 14(4):699–701, October 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-4-699.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        4-699.pdf.

Hussey:1936:OES

[HW36] L. W. Hussey and L. R. Wrathall. Oscillations in an electromechanical system. The Bell System Technical Journal, 15(3):441–445, July 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-3-441.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        3-441.pdf.

Ives:1930:ITS

[IGB30] Herbert E. Ives, Frank Gray, and M. W. Baldwin. Image transmission system for two-way television. The Bell System Technical Journal, 9(3):448–469, July 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        3-448.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-3-448.pdf.

Inglis:1932:VET

[IGJ32] A. H. Inglis, C. H. G. Gray, and R. T. Jenkins. A voice and ear for telephone measurements. The Bell System Technical Journal, 11(2):293–317, April 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-2-293.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
2-293.pdf.

Ilgenfritz:1939:LPD

[IHW39] L. M. Ilgenfritz, R. N. Hunter, and A. L. Whitman. Line problems in the development of the twelve-channel open-wire carrier system. The Bell System Technical Journal, 18(2):363–387, April 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-2-363.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
2-363.pdf.

Inglis:1938:TFN

[Ing38] A. H. Inglis. Transmission features of the new telephone sets. The Bell System Technical Journal, 17(3):358–380, July 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-3-358.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
3-358.pdf.

Ives:1931:MCT

[Ive31a] Herbert E. Ives. A multi-channel television apparatus. The Bell System Technical Journal, 10 (1):33–45, January 1931. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-1-33.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol10-
                                        1931/articles/bstj10-1-33.pdf.

Ives:1931:SOF

[Ive31b] Herbert E. Ives. Some optical features in two-way television. The Bell System Technical Journal, 10(2):265–272, April 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-2-265.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        2-265.pdf.

Jewett:1935:DGC

[Jew35] F. B. Jewett. Dr. George A. Campbell. The Bell System Technical Journal, 14(4):553–557, October 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-4-553.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        4-553.pdf.

Jones:1932:DHT

[JI32] W. C. Jones and A. H. Inglis. The development of a handset for telephone stations. The Bell System Technical Journal, 11 (2):245–263, April 1932. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-2-245.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
2-245.pdf.

Johnson:1935:LA

[JL35] J. B. Johnson and F. B. Llewellyn.

Limits to amplification. The Bell System Technical Journal, 14 (1):85–96, January 1935. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-1-85.pdf;   http://www.
alcatel-lucent.com/bstj/vol14-
1935/articles/bstj14-1-85.pdf.

Johnson:1932:CRO

[Joh32] J. B. Johnson. The cathode ray oscillograph. The Bell System Technical Journal, 11(1):1–27, January 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol11-
1932/articles/bstj11-1-1.pdf.

Jones:1931:NSS

[Jon31a] R. L. Jones. New standard specifications for wood poles. The Bell System Technical Journal, 10(3):514–524, July 1931. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-3-514.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        3-514.pdf.

Jones:1931:CCM

[Jon31b] W. C. Jones. Condenser and carbon microphones — their construction and use. The Bell System Technical Journal, 10 (1):46–62, January 1931. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-1-46.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol10-
                                        1931/articles/bstj10-1-46.pdf.

Jones:1938:INT

[Jon38] W. C. Jones. Instruments for the new telephone sets. The Bell System Technical Journal, 17(3):338–357, July 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-3-338.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        3-338.pdf.

Johnson:1939:SCM

[JS39] A. G. Johnson and L. I. Shaw. Some ceramic manufacturing developments of the western electric company. The Bell System Technical Journal, 18(2):255–279, April 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol18/
bstj18-2-255.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
2-255.pdf.

Kendall:1939:TCC

[KA39] B. W. Kendall and H. A. Affel. A twelve-channel carrier telephone system for open-wire lines. The Bell System Technical Journal, 18(1):119–142, January 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-1-119.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
1-119.pdf.

Kalb:1935:FDT

[KB35] Robert M. Kalb and William R.

Bennett. Ferromagnetic distortion of a two-frequency wave. The Bell System Technical Journal, 14(2):322–359, April 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-2-322.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
2-322.pdf.

Kemp:1931:PNI

[Kem31] A. R. Kemp. Paragutta, A new insulating material for submarine cables. The Bell System Technical Journal, 10(1):132–148, January 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-1-132.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        1-132.pdf.

Knowlton:1936:SSF

[KLS36] A. D. Knowlton, G. A. Locke, and F. J. Singer. Switchboards and signaling facilities of the teletypewriter exchange system. The Bell System Technical Journal, 15 (4):504–528, October 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-4-504.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        4-504.pdf.

Kelly:1935:VTH

[KS35] M. J. Kelly and A. L. Samuel. Vacuum tubes as high-frequency oscillators. The Bell System Technical Journal, 14(1):97–134, January 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-1-97.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol14-
                                        1935/articles/bstj14-1-97.pdf.

Lane:1930:PDT

[Lan30] C. E. Lane. Phase distortion in telephone apparatus. The Bell System Technical Journal, 9(3):493–521, July 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
3-493.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-3-493.pdf.

Lane:1938:CCF

[Lan38] C. E. Lane. Crystal channel filters for the cable carrier system. The Bell System Technical Journal, 17(1):125–136, January 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-1-125.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
1-125.pdf.

Lawther:1935:ANT

[Law35] H. P. Lawther, Jr. An application of number theory to the splicing of telephone cables. The Bell System Technical Journal, 14(2):273–284, April 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-2-273.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
2-273.pdf.

Llewellyn:1939:PUH

[LB39] F. B. Llewellyn and A. E. Bowen.

The production of ultra-high-frequency oscillations by means of diodes. The Bell System Technical Journal, 18(2):280–291, April 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/
                                        bstj18-2-280.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol18-1939/articles/bstj18-
                                        2-280.pdf.

Legg:1936:MML

[Leg36] Victor E. Legg. Magnetic measurements at low flux densities using the alternating current bridge. The Bell System Technical Journal, 15(1):39–62, January 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-1-39.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol15-
                                        1936/articles/bstj15-1-39.pdf.

Legg:1939:SMM

[Leg39] V. E. Legg. Survey of magnetic materials and applications in the telephone system. The Bell System Technical Journal, 18(3):438–464, July 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/
                                        bstj18-3-438.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol18-1939/articles/bstj18-
                                        3-438.pdf.

Llewellyn:1932:CFO

[Lle32] F. B. Llewellyn. Constant frequency oscillators. The Bell System Technical Journal, 11(1): 67–100, January 1932. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-1-67.pdf;   http://www.
alcatel-lucent.com/bstj/vol11-
1932/articles/bstj11-1-67.pdf.

Llewellyn:1934:VTE

[Lle34] F. B. Llewellyn. Vacuum tube electronics at ultra-high frequencies. The Bell System Technical Journal, 13(1):59–101, January 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-1-59.pdf;   http://www.
alcatel-lucent.com/bstj/vol13-
1934/articles/bstj13-1-59.pdf.

Llewellyn:1935:OUH

[Lle35]  F. B. Llewellyn.  Operation

of ultra-high-frequency vacuum tubes. The Bell System Technical Journal, 14(4):632–665, October 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-4-632.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
4-632.pdf.

Llewellyn:1936:ENN

[Lle36] F. B. Llewellyn. Equivalent networks of negative-grid vacuum tubes at ultra-high frequencies. The Bell System Technical Journal, 15(4):575–586, October 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/

                                        bstj15-4-575.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        4-575.pdf.

Levin:1934:ETT

[LP34] S. A. Levin and Liss C. Peterson. An extension of the theory of three-electrode vacuum tube circuits. The Bell System Technical Journal, 13(4):523–531, October 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol13/
                                        bstj13-4-523.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol13-1934/articles/bstj13-
                                        4-523.pdf.

Levin:1935:FET

[LP35] S. A. Levin and Liss C. Peterson. Further extensions of the theory of multi-electrode vacuum tube circuits. The Bell System Technical Journal, 14(4):666–679, October 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-4-666.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        4-666.pdf.

Lucas:1930:SNT

[Luc30] Francis F. Lucas. Structure and nature of troostite. The Bell System Technical Journal, 9(1): 101–120, January 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.

com/BSTJ/images/Vol09/bstj9-
1-101.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-1-101.pdf.

Lack:1934:SIQ

[LWF34] F. R. Lack, G. W. Willard, and I. E. Fair. Some improvements in quartz crystal circuit elements. The Bell System Technical Journal, 13(3):453–463, July 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-3-453.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
3-453.pdf.

MacNair:1930:ORT

[Mac30] Walter A. MacNair. Optimum reverberation time for auditoriums. The Bell System Technical Journal, 9(2):390–397, April 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
2-390.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-2-390.pdf.

Martin:1930:TFR

[Mar30] W. H. Martin. Transmitted frequency range for telephone message circuits. The Bell System Technical Journal, 9(3):483–486, July 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-

                                        3-483.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-3-483.pdf.

Martin:1931:RTP

[Mar31] W. H. Martin. Rating the transmission performance of telephone circuits. The Bell System Technical Journal, 10(1): 116–131, January 1931. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-1-116.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        1-116.pdf.

Mason:1930:ANA

[Mas30] W. P. Mason. The approximate networks of acoustic filters. The Bell System Technical Journal, 9 (2):332–340, April 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        2-332.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-2-332.pdf.

Mason:1934:EWF

[Mas34] W. P. Mason. Electrical wave filters employing quartz crystals as elements. The Bell System Technical Journal, 13(3):405–452, July 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol13/
                                        bstj13-3-405.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol13-1934/articles/bstj13-
3-405.pdf.

Mason:1935:TDE

[Mas35] W. P. Mason. Technical digests: An electromechanical representation of a piezoelectric crystal used as A transducer. The Bell System Technical Journal, 14 (4):718–723, October 1935. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-4-718.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
4-718.pdf.

Mason:1937:RCB

[Mas37] W. P. Mason. Resistance compensated band-pass crystal filters for use in unbalanced circuits. The Bell System Technical Journal, 16(4):423–436, October 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-4-423.pdf;   http://
www.alcatel-lucent.com/bstj/
vol16-1937/articles/bstj16-
4-423.pdf.

Martin:1930:TDT

[MD30] W. H. Martin and W. F. Davidson. The trend in the design of telephone transmitters and receivers. The Bell System Technical Journal, 9(4): 622–627, October 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        4-622.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-4-622.pdf.

McKown:1933:SET

[ME33] F. W. McKown and J. W. Emling. A system of effective transmission data for rating telephone circuits. The Bell System Technical Journal, 12(3):331–346, July 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol12/
                                        bstj12-3-331.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol12-1933/articles/bstj12-
                                        3-331.pdf.

Meacham:1938:BSO

[Mea38] L. A. Meacham. The bridge stabilized oscillator. The Bell System Technical Journal, 17(4): 574–591, October 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-4-574.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        4-574.pdf.

Mertz:1934:TSR

[MG34] Pierre Mertz and Frank Gray. A theory of scanning and its relation to the characteristics of the transmitted signal in telephotography and television. The Bell System Technical Journal, 13(3):464–515, July 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-3-464.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
3-464.pdf.

Martin:1931:SCW

[MH31] J. C. Martin and H. L. Huber. Status of cooperative work on joint use of poles. The Bell System Technical Journal, 10 (2):231–240, April 1931. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-2-231.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
2-231.pdf.

Murphy:1937:DPI

[MM37] E. J. Murphy and S. O. Morgan. The dielectric properties of insulating materials. The Bell System Technical Journal, 16(4): 493–512, October 1937. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-4-493.pdf;   http://
www.alcatel-lucent.com/bstj/
vol16-1937/articles/bstj16-
4-493.pdf.

Murphy:1938:DPI

[MM38] E. J. Murphy and S. O. Morgan. The dielectric properties of insulating materials. The Bell System Technical Journal, 17(4): 640–669, October 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-4-640.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        4-640.pdf.

Murphy:1939:DPI

[MM39] E. J. Murphy and S. O. Morgan. The dielectric properties of insulating materials, III alternating and direct current conductivity. The Bell System Technical Journal, 18(3):502–537, July 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/
                                        bstj18-3-502.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol18-1939/articles/bstj18-
                                        3-502.pdf.

Molina:1931:BTE

[Mol31] Edward C. Molina. Bayes’ theorem — an expository presentation. The Bell System Technical Journal, 10(2):273–283, April 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-2-273.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        2-273.pdf.

Molina:1932:ELI

[Mol32] Edward C. Molina. An expansion for Laplacian integrals in terms of incomplete gamma functions, and some applications. The Bell System Technical Journal, 11(4):563–575, October 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-4-563.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
4-563.pdf.

Molina:1936:LEH

[Mol36] E. C. Molina. A Laplacian expansion for Hermitian–Laplace functions of high order. The Bell System Technical Journal, 15(3):355–362, July 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-3-355.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
3-355.pdf.

Montgomery:1938:OHA

[Mon38] H. C. Montgomery. An optical harmonic analyzer. The Bell System Technical Journal, 17(3):406–415, July 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-3-406.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
3-406.pdf.

Mertz:1937:IBB

[MP37] Pierre Mertz and K. W. Pfleger. Irregularities in broad-band wire transmission circuits. The Bell System Technical Journal, 16(4): 541–559, October 1937. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-4-541.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        4-541.pdf.

McLean:1932:SPP

[MPS32] D. A. McLean, R. L. Peek, Jr., and E. E. Schumacher. Some physical properties of wiping solders. The Bell System Technical Journal, 11(1):101–125, January 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol11/
                                        bstj11-1-101.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol11-1932/articles/bstj11-
                                        1-101.pdf.

Marshall:1936:NDM

[MR36] R. N. Marshall and F. F. Romanow. A non-directional microphone. The Bell System Technical Journal, 15(3):405–423, July 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-3-405.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        3-405.pdf.

Mason:1937:UCB

[MS37] W. P. Mason and R. A. Sykes.

The use of coaxial and balanced transmission lines in filters and wide band transformers for high radio frequencies. The Bell System Technical Journal, 16(3):275–302, July 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-3-275.pdf;   http://
www.alcatel-lucent.com/bstj/
vol16-1937/articles/bstj16-
3-275.pdf.

Mathes:1934:CAA

[MW34] R. C. Mathes and S. B. Wright.

The compandor — an aid against static in radio telephony. The Bell System Technical Journal, 13(3):315–332, July 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-3-315.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
3-315.pdf.

Nance:1932:WCA

[Nan32] H. H. Nance. Wire communication aids to air transportation. The Bell System Technical Journal, 11(3):462–476, July 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-3-462.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol11-1932/articles/bstj11-
                                        3-462.pdf.

Nyquist:1930:MPD

[NB30] H. Nyquist and S. Brand. Measurement of phase distortion. The Bell System Technical Journal, 9(3):522–549, July 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        3-522.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-3-522.pdf.

Nelson:1930:RBT

[Nel30] Edward L. Nelson. Radio broadcasting transmitters and related transmission phenomena. The Bell System Technical Journal, 9 (1):121–140, January 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        1-121.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-1-121.pdf.

Norwine:1938:CTI

[NM38] A. C. Norwine and O. J. Murphy. Characteristic time intervals in telephonic conversation. The Bell System Technical Journal, 17(2):281–291, April 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-2-281.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol17-1938/articles/bstj17-
2-281.pdf.

Norton:1937:CRA

[Nor37] E. L. Norton. Constant resistance with applications to filter groups. The Bell System Technical Journal, 16(2):178–193, April 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-2-178.pdf;   http://
www.alcatel-lucent.com/bstj/
vol16-1937/articles/bstj16-
2-178.pdf.

Norwine:1938:DCA

[Nor38] A. C. Norwine. Devices for controlling amplitude characteristics of telephonic signals. The Bell System Technical Journal, 17(4): 539–554, October 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-4-539.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
4-539.pdf.

Nyquist:1932:RT

[Nyq32] H. Nyquist. Regeneration theory. The Bell System Technical Journal, 11(1):126–147, January 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-1-126.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol11-1932/articles/bstj11-
                                        1-126.pdf.

OLeary:1939:ITC

[OBB39] J. T. O’Leary, E. C. Blessing, and J. W. Beyer. An improved three-channel carrier telephone system. The Bell System Technical Journal, 18(1):49–75, January 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/
                                        bstj18-1-49.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol18-
                                        1939/articles/bstj18-1-49.pdf.

Osborne:1932:EDJ

[OD32] H. S. Osborne and A. M. Dowling. The electrical discoveries of Joseph Henry, A supplement to the Bell System Technical Journal, July 1932. The Bell System Technical Journal, 11(3):1–22, July 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol11/
                                        bstj11-3-1.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol11-
                                        1932/articles/bstj11-3-1.pdf.

Osbourne:1930:GSP

[Osb30] H. S. Osbourne. A general switching plan for telephone toll service. The Bell System Technical Journal, 9(3):429–447, July 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        3-429.pdf;   http://www.alcatel-

lucent.com/bstj/vol09-1930/
articles/bstj9-3-429.pdf.

Oswald:1930:TTS

[Osw30] A. A. Oswald. Transoceanic telephone service — short-wave equipment. The Bell System Technical Journal, 9(2): 270–289, April 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
2-270.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-2-270.pdf.

Pack:1931:SCP

[Pac31] R. F. Pack. Symposium on coordination of power and telephone plant. The Bell System Technical Journal, 10(2):155–158, April 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-2-155.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
2-155.pdf.

Payne:1930:ICW

[Pay30] E. B. Payne. Impedance correction of wave filters. The Bell System Technical Journal, 9(4): 770–793, October 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
4-770.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-4-770.pdf.

Pierce:1936:TST

[PB36] R. E. Pierce and E. W. Bemis. A transmission system for teletypewriter exchange service. The Bell System Technical Journal, 15(4):529–548, October 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-4-529.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
4-529.pdf.

Pearson:1934:FNV

[Pea34] G. L. Pearson. Fluctuation noise in vacuum tubes. The Bell System Technical Journal, 13(4): 634–653, October 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-4-634.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
4-634.pdf.

Penick:1934:MRM

[Pen34] D. B. Penick. The measurement and reduction of microphonic noise in vacuum tubes. The Bell System Technical Journal, 13(4):614–633, October 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-4-614.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
4-614.pdf.

Peterson:1930:TPG

[Pet30] Liss C. Peterson. Transients in parallel grounded circuits, one of which is of infinite length. The Bell System Technical Journal, 9 (4):760–769, October 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
4-760.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-4-760.pdf.

Peterson:1939:IPE

[Pet39] Liss C. Peterson. Impedance properties of electron streams. The Bell System Technical Journal, 18(3):465–481, July 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-3-465.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
3-465.pdf.

Peterson:1939:EMC

[PH39] E. Peterson and L. W. Hussey.

Equivalent modulator circuits. The Bell System Technical Journal, 18(1):32–48, January 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-1-32.pdf;   http://www.
alcatel-lucent.com/bstj/vol18-
1939/articles/bstj18-1-32.pdf.

Pidgeon:1935:TME

[Pid35] H. A. Pidgeon. Theory of multielectrode vacuum tubes. The Bell System Technical Journal, 14 (1):44–84, January 1935. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-1-44.pdf;   http://www.
alcatel-lucent.com/bstj/vol14-
1935/articles/bstj14-1-44.pdf.

Pilliod:1939:TTL

[Pil39] J. J. Pilliod. Transcontinental telephone lines. The Bell System Technical Journal, 18(1): 235–245, January 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-1-235.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
1-235.pdf.

Prescott:1932:COS

[PK32] C. J. Prescott, Jr. and M. J. Kelly. The caesium–oxygen– silver photoelectric cell. The Bell System Technical Journal, 11(3):334–367, July 1932. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-3-334.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
3-334.pdf.

Peterson:1934:RTE

[PKW34] E. Peterson, J. G. Kreer, and L. A. Ware. Regeneration theory and experiment. The Bell System Technical Journal, 13(4): 680–700, October 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-4-680.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
4-680.pdf.

Pennell:1936:MES

[PL36] W. O. Pennell and H. P. Lawther.

A magneto-elastic source of noise in steel telephone wires. The Bell System Technical Journal, 15 (2):334–339, April 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-2-334.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
2-334.pdf.

Peterson:1937:MGG

[PMW37] E. Peterson, J. M. Manley, and L. R. Wrathall. Magnetic generation of a group of harmonics. The Bell System Technical Journal, 16(4):437–455, October 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-4-437.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol16-1937/articles/bstj16-
                                        4-437.pdf.

Potter:1936:RSW

[PP36] R. K. Potter and A. C. Peterson, Jr. The reliability of short-wave radio telephone circuits. The Bell System Technical Journal, 15(2):181–196, April 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-2-181.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        2-181.pdf.

Polkinghorn:1935:SSS

[PS35] F. A. Polkinghorn and N. F. Schlaack. A single-sideband shortwave system for transatlantic telephony. The Bell System Technical Journal, 14(3):489–508, July 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-3-489.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        3-489.pdf.

Quarles:1930:MTW

[Qua30] D. A. Quarles. Motion of telephone wires in wind. The Bell System Technical Journal, 9 (2):356–363, April 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        2-356.pdf;   http://www.alcatel-

lucent.com/bstj/vol09-1930/
articles/bstj9-2-356.pdf.

Quarles:1936:NTU

[Qua36] D. A. Quarles. A new type of underground telephone wire. The Bell System Technical Journal, 15(3):446–454, July 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-3-446.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
3-446.pdf.

Rack:1938:ESC

[Rac38] A. J. Rack. Effect of space charge and transit time on the shot noise in diodes. The Bell System Technical Journal, 17(4):592–619, October 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-4-592.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
4-592.pdf.

Reynolds:1936:NTS

[Rey36] F. W. Reynolds. A new telephotograph system. The Bell System Technical Journal, 15(4): 549–574, October 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-4-549.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol15-1936/articles/bstj15-
                                        4-549.pdf.

Rice:1937:SWF

[Ric37] S. O. Rice. Series for the wave function of a radiating dipole at the Earth’s surface. The Bell System Technical Journal, 16(1): 101–109, January 1937. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-1-101.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        1-101.pdf.

Riordan:1931:TGW

[Rio31] John Riordan. Transients in grounded wires on the Earth’s surface. The Bell System Technical Journal, 10(3):420–431, July 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-3-420.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        3-420.pdf.

Riordan:1937:LNT

[Rio37] John Riordan. A ladder network theorem. The Bell System Technical Journal, 16(3):303–318, July 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-3-303.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol16-1937/articles/bstj16-
3-303.pdf.

Riordan:1939:NIT

[Rio39] John Riordan. The number of impedances of an n terminal network. The Bell System Technical Journal, 18(2):300–314, April 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-2-300.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
2-300.pdf.

Riordan:1933:MIG

[RS33] John Riordan and Erling D.

Sunde. Mutual impedance of grounded wires for horizontally stratified two-layer Earth. The Bell System Technical Journal, 12 (2):162–177, April 1933. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-2-162.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
2-162.pdf.

Samuel:1937:TDN

[Sam37] A. L. Samuel. Technical digests — A negative-grid triode oscillator and amplifier for ultra-high frequencies. The Bell System Technical Journal, 16(4):568–573, October 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/

                                        bstj16-4-568.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        4-568.pdf.

Schelleng:1933:USW

[SBF33] J. C. Schelleng, C. R. Burrows, and E. B. Ferrell. Ultrashort wave propagation. The Bell System Technical Journal, 12 (2):125–161, April 1933. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol12/
                                        bstj12-2-125.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol12-1933/articles/bstj12-
                                        2-125.pdf.

Shanck:1939:RDM

[SCC39] R. B. Shanck, F. A. Cowan, and S. I. Cory. Recent developments in the measurement of telegraph transmission. The Bell System Technical Journal, 18(1): 143–189, January 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/
                                        bstj18-1-143.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol18-1939/articles/bstj18-
                                        1-143.pdf.

Schelkunoff:1934:ETC

[Sch34] S. A. Schelkunoff. The electromagnetic theory of coaxial transmission lines and cylindrical shields. The Bell System Technical Journal, 13(4):532–579, October 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol13/
bstj13-4-532.pdf;   http://
www.alcatel-lucent.com/bstj/
vol13-1934/articles/bstj13-
4-532.pdf.

Schelkunoff:1936:SET

[Sch36] S. A. Schelkunoff. Some equivalence theorems of electromagnetics and their application to radiation problems. The Bell System Technical Journal, 15(1): 92–112, January 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-1-92.pdf;   http://www.
alcatel-lucent.com/bstj/vol15-
1936/articles/bstj15-1-92.pdf.

Schelkunoff:1938:ICA

[Sch38] S. A. Schelkunoff. The impedance concept and its application to problems of reflection, refraction, shielding and power absorption. The Bell System Technical Journal, 17(1):17–48, January 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-1-17.pdf;   http://www.
alcatel-lucent.com/bstj/vol17-
1938/articles/bstj17-1-17.pdf.

Scott:1932:MSP

[Sco32] K. L. Scott. Magnet steels and permanent magnets — relationships among their magnetic properties. The Bell System Technical Journal, 11(3):383–401, July 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol11/
                                        bstj11-3-383.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol11-1932/articles/bstj11-
                                        3-383.pdf.

Scriven:1934:RAP

[Scr34] E. O. Scriven. Reproduction in auditory perspective. The Bell System Technical Journal, 13 (2):278–284, April 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol13/
                                        bstj13-2-278.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol13-1934/articles/bstj13-
                                        2-278.pdf.

Sterba:1932:TLS

[SF32] E. J. Sterba and C. B. Feldman. Transmission lines for short-wave radio systems. The Bell System Technical Journal, 11(3):411–450, July 1932. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol11/
                                        bstj11-3-411.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol11-1932/articles/bstj11-
                                        3-411.pdf.

Schafer:1936:TDE

[SG36] J. P. Schafer and W. M. Goodall. Technical digests—eclipse effects in the ionsphere. The Bell System Technical Journal, 15(1): 162–166, January 1936. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-1-162.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
1-162.pdf.

Shewhart:1930:EQC

[She30] W. A. Shewhart. Economic quality control of manufactured product. The Bell System Technical Journal, 9(2):364–389, April 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
2-364.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-2-364.pdf.

Shea:1931:DML

[She31] John R. Shea. Developments in the manufacture of lead-covered paper-insulated telephone cable. The Bell System Technical Journal, 10(3):432–471, July 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-3-432.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
3-432.pdf.

Shockley:1939:QPS

[Sho39] W. Shockley. The quantum physics of solids, I. the energies of electrons in crystals. The Bell System Technical Journal, 18(4): 645–723, October 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/
                                        bstj18-4-645.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol18-1939/articles/bstj18-
                                        4-645.pdf.

Sivian:1931:ACC

[Siv31] L. J. Sivian. Absolute calibration of condenser transmitters. The Bell System Technical Journal, 10(1):96–115, January 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-1-96.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol10-
                                        1931/articles/bstj10-1-96.pdf.

Skellett:1936:TDC

[Ske36] A. M. Skellett. Technical digests— on the correlation of radio transmission with solar phenomena. The Bell System Technical Journal, 15(1):157–161, January 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-1-157.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        1-157.pdf.

Spain:1936:TDR

[SLT36] C. J. Spain, D. P. Loye, and E. W. Templin. Technical digests— reduction of airplane noise and vibration. The Bell System Technical Journal, 15(4):626–627, October 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol15/
bstj15-4-626.pdf;   http://
www.alcatel-lucent.com/bstj/
vol15-1936/articles/bstj15-
4-626.pdf.

Starbird:1939:SAT

[SM39] L. C. Starbird and J. D. Mathis.

Some applications of the type ‘J’ carrier system. The Bell System Technical Journal, 18 (2):338–362, April 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-2-338.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
2-338.pdf.

Snow:1931:AFR

[Sno31] W. B. Snow. Audible frequency ranges of music, speech and noise. The Bell System Technical Journal, 10(4):616–627, October 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-4-616.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
4-616.pdf.

Schelkunoff:1937:CBC

[SO37] S. A. Schelkunoff and T. M.

Odarenko. Crosstalk between coaxial transmission lines. The Bell System Technical Journal, 16 (2):144–164, April 1937. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-2-144.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        2-144.pdf.

Southworth:1931:CFA

[Sou31] G. C. Southworth. Certain factors affecting the gain of directive antennas. The Bell System Technical Journal, 10(1):63–95, January 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-1-63.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol10-
                                        1931/articles/bstj10-1-63.pdf.

Southworth:1936:HFW

[Sou36] G. C. Southworth. Hyperfrequency wave guides—general considerations and experimental results. The Bell System Technical Journal, 15(2):284–309, April 1936. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol15/
                                        bstj15-2-284.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol15-1936/articles/bstj15-
                                        2-284.pdf.

Schlaack:1935:UUS

[SP35] N. F. Schlaack and F. A. Polkinghorn. An unattended ultrashort-wave radio telephone system. The Bell System Technical Journal, 14(3):534–541, July 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-3-534.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
3-534.pdf.

Scudder:1939:CDT

[SR39] F. J. Scudder and J. N. Reynolds.

Crossbar dial telephone switching system. The Bell System Technical Journal, 18(1):76–118, January 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-1-76.pdf;   http://www.
alcatel-lucent.com/bstj/vol18-
1939/articles/bstj18-1-76.pdf.

Scaff:1933:STP

[SS33] J. H. Scaff and E. E. Schumacher. Some theoretical and practical aspects of gases in metals. The Bell System Technical Journal, 12(2):178–196, April 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-2-178.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
2-178.pdf.

Steinberg:1934:PFS

[SS34] J. C. Steinberg and W. B. Snow.

Physical factors, symposium on wire transmission of symphonic music and its reproduction in auditory perspective. The Bell System Technical Journal, 13 (2):245–258, April 1934. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol13/
                                        bstj13-2-245.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol13-1934/articles/bstj13-
                                        2-245.pdf.

Samuel:1937:PAU

[SS37] A. L. Samuel and N. E. Sowers. A power amplifier for ultra-high frequencies. The Bell System Technical Journal, 16(1):10–34, January 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-1-10.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol16-
                                        1937/articles/bstj16-1-10.pdf.

Steinberg:1930:EPD

[Ste30] John C. Steinberg. Effects of phase distortion on telephone quality. The Bell System Technical Journal, 9(3):550–566, July 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.
                                        com/BSTJ/images/Vol09/bstj9-
                                        3-550.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol09-1930/
                                        articles/bstj9-3-550.pdf.

Stevenson:1938:SFO

[Ste38]  G. H. Stevenson.  Stabilized

feedback oscillators. The Bell System Technical Journal, 17 (3):458–474, July 1938. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-3-458.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
3-458.pdf.

Stoller:1930:SST

[Sto30] H. M. Stoller. Synchronization system for two-way television. The Bell System Technical Journal, 9(3):470–477, July 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
3-470.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-3-470.pdf.

Strieby:1937:MCT

[Str37] M. E. Strieby. A million-cycle telephone system. The Bell System Technical Journal, 16 (1):1–9, January 1937. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol16-
1937/articles/bstj16-1-1.pdf.

Strieby:1938:CCS

[Str38] M. E. Strieby. Coaxial cable system for television transmission. The Bell System Technical Journal, 17(3):438–457, July 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-3-438.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        3-438.pdf.

Sunde:1937:TDC

[Sun37] E. D. Sunde. Technical digests — currents and potentials along leaky ground-return conductors. The Bell System Technical Journal, 16(1):110–112, January 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-1-110.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        1-110.pdf.

Taylor:1937:RTN

[Tay37] C. C. Taylor. Radio telephone noise reduction by voice control at receiver. The Bell System Technical Journal, 16(4):475–486, October 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-4-475.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        4-475.pdf.

Townsend:1939:PMT

[TC39] J. R. Townsend and W. J. Clarke. Plastic materials in telephone use. The Bell System Technical Journal, 18(3):482–501, July 1939. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol18/
bstj18-3-482.pdf;   http://
www.alcatel-lucent.com/bstj/
vol18-1939/articles/bstj18-
3-482.pdf.

Thuras:1935:EFG

[TJO35] A. L. Thuras, R. T. Jenkins, and H. T. O’Neil. Extraneous frequencies generated in air carrying intense sound waves. The Bell System Technical Journal, 14 (1):159–172, January 1935. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-1-159.pdf;   http://
www.alcatel-lucent.com/bstj/
vol14-1935/articles/bstj14-
1-159.pdf.

Walker:1933:EAH

[Wal33] Albert C. Walker. Effect of atmospheric humidity and temperature on the relation between moisture content and electrical conductivity of cotton. The Bell System Technical Journal, 12(4):431–451, October 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol12/
bstj12-4-431.pdf;   http://
www.alcatel-lucent.com/bstj/
vol12-1933/articles/bstj12-
4-431.pdf.

Walker:1937:MT

[Wal37] Albert C. Walker. Moisture in textiles. The Bell System Technical Journal, 16(2):228–246, April 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-2-228.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol16-1937/articles/bstj16-
                                        2-228.pdf.

Watson:1938:FTU

[Wat38] E. F. Watson. Fundamentals of teletypewriters used in the Bell System. The Bell System Technical Journal, 17(4):620–639, October 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-4-620.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        4-620.pdf.

Wills:1931:SJD

[WB31] H. L. Wills and O. B. Blackwell. Status of joint development and research on noise frequency induction. The Bell System Technical Journal, 10(2):184–205, April 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-2-184.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        2-184.pdf.

Wood:1932:CAT

[WB32] E. B. Wood and D. R. Brobst. Cellulose acetate treatment for textile insulation — engineering development. The Bell System Technical Journal, 11(2):213–230, April 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-2-213.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
2-213.pdf.

Wilson:1930:RTS

[WE30] William Wilson and Lloyd Espenschied. Radio telephone service to ships at sea. The Bell System Technical Journal, 9(3):407–428, July 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
3-407.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-3-407.pdf.

Wegel:1930:TVL

[Weg30] R. L. Wegel. Theory of vibration of the larynx. The Bell System Technical Journal, 9(1): 207–227, January 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
1-207.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-1-207.pdf.

Wente:1935:AI

[Wen35] E. C. Wente. Acoustical instruments. The Bell System Technical Journal, 14(3):388–412, July 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol14/
                                        bstj14-3-388.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        3-388.pdf.

Walker:1933:PIT

[WF33] H. G. Walker and L. S. Ford. Pulp insulation for telephone cables. The Bell System Technical Journal, 12(1):1–21, January 1933. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol12/
                                        bstj12-1-1.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol12-
                                        1933/articles/bstj12-1-1.pdf.

Wiebusch:1939:DCS

[Wie39] C. F. Wiebusch. Dial clutch of the spring type. The Bell System Technical Journal, 18(4): 724–741, October 1939. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol18/
                                        bstj18-4-724.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol18-1939/articles/bstj18-
                                        4-724.pdf.

Williams:1930:CTI

[Wil30a] Robert R. Williams. Chemistry in the telephone industry. The Bell System Technical Journal, 9 (4):603–621, October 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

                                        URL http://bstj.bell-labs.

com/BSTJ/images/Vol09/bstj9-
4-603.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-4-603.pdf.

Wilson:1930:STL

[Wil30b] L. T. Wilson. A study of telephone line insulators. The Bell System Technical Journal, 9(4): 697–729, October 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
4-697.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-4-697.pdf.

Wilkinson:1931:ITS

[Wil31] R. I. Wilkinson. The interconnection of telephone systems — graded multiples. The Bell System Technical Journal, 10(4): 531–564, October 1931. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-4-531.pdf;   http://
www.alcatel-lucent.com/bstj/
vol10-1931/articles/bstj10-
4-531.pdf.

Willets:1935:TDS

[Wil35] H. N. Willets. Technical digests:

Ship sets for harbor ship-to-shore service. The Bell System Technical Journal, 14(4):713–717, October 1935. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol14/
bstj14-4-713.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol14-1935/articles/bstj14-
                                        4-713.pdf.

Wise:1931:EGP

[Wis31] W. Howard Wise. Effect of ground permeability on ground return circuits. The Bell System Technical Journal, 10(3):472–484, July 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-3-472.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        3-472.pdf.

Wise:1937:PRZ

[Wis37] W. Howard Wise. The physical reality of Zenneck’s surface wave. The Bell System Technical Journal, 16(1):35–44, January 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-1-35.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol16-
                                        1937/articles/bstj16-1-35.pdf.

Waterman:1937:LEW

[WLH37] R. E. Waterman, John Leutritz, and Caleb M. Hill. A laboratory evaluation of wood preservatives. The Bell System Technical Journal, 16(2):194–211, April 1937. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol16/
                                        bstj16-2-194.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol16-1937/articles/bstj16-
2-194.pdf.

Williams:1930:SRN

[WM30] W. J. Williams and Ralph G. Mc-Curdy. A survey of room noise in telephone locations. The Bell System Technical Journal, 9(4): 652–667, October 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
4-652.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-4-652.pdf.

Wright:1932:TWR

[WM32] S. B. Wright and D. Mitchell.

Two-way radio telephone circuits. The Bell System Technical Journal, 11(3):368–382, July 1932. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol11/
bstj11-3-368.pdf;   http://
www.alcatel-lucent.com/bstj/
vol11-1932/articles/bstj11-
3-368.pdf.

Wright:1937:V

[Wri37] S. B. Wright. The vodas. The Bell System Technical Journal, 16 (4):456–474, October 1937. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol16/
bstj16-4-456.pdf;   http://
www.alcatel-lucent.com/bstj/
vol16-1937/articles/bstj16-
4-456.pdf.

Wright:1938:ARC

[Wri38] S. B. Wright. Amplitude range control. The Bell System Technical Journal, 17(4):520–538, October 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol17/
                                        bstj17-4-520.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol17-1938/articles/bstj17-
                                        4-520.pdf.

Wente:1931:MCT

[WT31] E. C. Wente and A. L. Thuras. Moving-coil telephone receivers and microphones. The Bell System Technical Journal, 10(4): 565–576, October 1931. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol10/
                                        bstj10-4-565.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol10-1931/articles/bstj10-
                                        4-565.pdf.

Wente:1934:LSM

[WT34] E. C. Wente and A. L. Thuras. Loud speakers and microphones, symposium on wire transmission of symphonic music and its reproduction in auditory perspective. The Bell System Technical Journal, 13(2):259–277, April 1934. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol13/
                                        bstj13-2-259.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol13-1934/articles/bstj13-
2-259.pdf.

Weaver:1938:CNF

[WTD38] M. A. Weaver, R. S. Tucker, and P. S. Darnell. Crosstalk and noise features of cable carrier telephone system. The Bell System Technical Journal, 17(1):137–161, January 1938. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol17/
bstj17-1-137.pdf;   http://
www.alcatel-lucent.com/bstj/
vol17-1938/articles/bstj17-
1-137.pdf.

Zinn:1930:WPC

[Zin30] M. K. Zinn. Wave propagation over continuously loaded fine wires. The Bell System Technical Journal, 9(1): 189–206, January 1930. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic).

URL http://bstj.bell-labs.
com/BSTJ/images/Vol09/bstj9-
1-189.pdf;   http://www.alcatel-
lucent.com/bstj/vol09-1930/
articles/bstj9-1-189.pdf.

Zobel:1931:ETD

[Zob31] Otto J. Zobel. Extensions to the theory and design of electric wavefilters. The Bell System Technical Journal, 10(2):284–341, April 1931. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol10/
bstj10-2-284.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol10-1931/articles/bstj10-
                                        2-284.pdf.